### v2

# COMMA: A Communicative Multimodal Multi-Agent Benchmark

Abstract
The rapid advances of multimodal agents built on large foundation models have largely overlooked their potential for language-based communication between agents in collaborative tasks. This oversight presents a critical gap in understanding their effectiveness in real-world deployments, particularly when communicating with humans. Existing agentic benchmarks fail to address key aspects of inter-agent communication and collaboration, particularly in scenarios where agents have unequal access to information and must work together to achieve tasks beyond the scope of individual capabilities. To fill this gap, we introduce a novel benchmark designed to evaluate the collaborative performance of multimodal multi-agent systems through language communication. Our benchmark features a variety of scenarios, providing a comprehensive evaluation across four key categories of agentic capability in a communicative collaboration setting. By testing both agent-agent and agent-human collaborations using open-source and closed-source models, our findings reveal surprising weaknesses in state-of-the-art models, including proprietary models like GPT-4o. Some of these models struggle to outperform even a simple random agent baseline in agent-agent collaboration and only surpass the random baseline when a human is involved.
Machine Learning, ICML
1 Introduction
The field of multimodal agents is experiencing rapid growth  (Xu et al., 2024b; Xie et al., 2024; Cao et al., 2024), with research efforts expanding at an unprecedented pace. However, amidst this growth, a critical gap in research has emerged: the lack of focus on collaborative work (Gurcan, 2024; Park et al., 2023; Hong et al., 2024; Liu et al., 2024b) among multiple multimodal agents. Synergistic operation of such agents is a highly promising but largely unexplored domain. Language agents can collaboratively finish complex tasks such as software development  (Qian et al., 2024; Du et al., 2024) or even machine learning research  (Schmidgall et al., 2025) by assuming functional roles such as system designer, function generator, etc. Current research on multimodal agents (Xu et al., 2024b; Xie et al., 2024; Cao et al., 2024) has mainly focused on individual agent capabilities, neglecting the potential for inter-agent collaboration. This limitation is further compounded by existing benchmarks such as TheAgentCompany (Xu et al., 2024a), VisualWebArena (Koh et al., 2024) and MME-RealWorld (Zhang et al., 2024), which do not assess collaborative performance between agents. As a result, our ability to evaluate and improve multi-agent systems remains constrained, hindering progress in this crucial area.
Several critical questions emerge in the context of multimodal agent collaboration. How can different agents effectively communicate multimodal information through language when they have varying levels of access to information? In scenarios where different agents possess diverse task-specific capabilities, how can they collaborate to accomplish objectives beyond the scope of any individual agent? These research settings remain largely uncharted and present significant challenges. Furthermore, the ability of agents to handle incomplete information is of paramount importance, particularly when working with sensitive data  (Li et al., 2024) (i.e. Agent application in healthcare where privacy concerns are critical  (Tang et al., 2024)). Exploration of these questions is crucial for advancing the field of multimodal agent collaboration. By addressing these challenges, we can expand the applicability of multimodal agents in real-world scenarios (Zhang et al., 2024), particularly those involving sensitive or restricted information.
Figure 1: Overview of the interaction between the Solver and Expert agents in our benchmark. Both agents operate with structured input corresponding to working and episodic memory. The Solver receives an image of the puzzle state (working memory) and makes decisions based on the available actions described in the task prompt. The Expert, guided by instruction manuals (working memory), provides advice based on the Solver’s descriptions, such as indicating which buttons to press. The Solver can choose to execute actions by interacting with the environment or communicate with the Expert for further guidance. Their interaction is documented through a dialogue, showcasing the cooperation required to complete the task. Both agents engage in self-reflection by referencing the conversation history, which is continuously updated and incorporated into their input as episodic memory.
Motivated by these aforementioned issues, we propose a novel benchmark for evaluating collaborative multimodal multi-agent frameworks to address critical gaps in current approaches (see  Figure 1). Our evaluation setting also simulates a scenario where an in-house agent with direct access to sensitive data (i.e., the AI solver) collaborates with external expert agents (i.e., the AI expert) to analyze information without compromising privacy. This evaluation setting simulates how we handle and extract insights from sensitive datasets across various domains.
We assess multimodal multi-agent systems using a series of carefully designed collaborative puzzle games. Building on cognitive science research employing simple puzzles to assess cognitive ability  (Davidson et al., 2006; St Clair-Thompson & Gathercole, 2006), we design vision-language puzzles to evaluate the core cognitive abilities of multimodal agents, including learning, adaptation, and problem-solving. These scenarios typically involve two-player setups where agents have access to different, complementary information. (i.e., in a bomb defusal game, one agent possesses details about the bomb, while the other has access to a disarming manual). By employing such diverse and interactive scenarios, we aim to provide a thorough assessment of multimodal multi-agent performance.
Our benchmark includes 10 distinct, easily customizable puzzles with thousands of unique solutions. We tested two different settings (AI-AI and AI-Human) and evaluated several popular multimodal models, including closed-source models (GPT-4V  (Achiam et al., 2023), GPT-4o  (Hurst et al., 2024)) and open-source models (QwenVL (Bai et al., 2023), and InternVL  (Chen et al., 2024)). Surprisingly, even powerful closed-source models such as the GPT series do not outperform a human solver, highlighting a potential growth area for future model development. Our contributions are as follows:
  • We propose an evaluation framework called COMMA, a multimodal agent benchmark focusing on language communication between multiple agents (Section 3).
  • Using COMMA, we carefully record conversations and performance metrics between state-of-the-art multimodal models such as QwenVL, InternVL, LLaMA, Gemini, GPT-4o, etc (Section 4).
  • We categorize the agent capabilities tested in our model and common failure modes, providing insight into potential future research directions for improving inter-agent communication (Section 5).
2 Related Work
Multi-agent Frameworks:
There are many emergent agent collaboration works  (Gurcan, 2024; Park et al., 2023; Hong et al., 2024; Liu et al., 2024b; Ghafarollahi & Buehler, 2024; Li et al., 2023; Wu et al., 2023) among multiple language agents. Multi-agent systems arise mainly in two different scenarios: (1) role-playing different task executors (e.g., software development requiring different roles of agents, such as program manager, software architect, programmer  (Du et al., 2024; Qian et al., 2024; Hong et al., 2024), scientific discovery simulation  (Wu et al., 2023), and social simulation  (Park et al., 2023; Gurcan, 2024; Park et al., 2024)); (2) communicating between agents with different pieces of information  (Wu et al., 2023; Li et al., 2023) (e.g., consulting experts without sharing some sensitive or confidential data. In our case, the AI solver has some private multimodal data, and the AI expert has domain-specific knowledge or instructions).
Instruction-based Agent Benchmarks:
Instruction-based agent benchmarks evaluate an agent’s capability of following a human instructions to finish a task (e.g., navigating on a website, interacting with an operating system  (Xu et al., 2024b; Xie et al., 2024; Cao et al., 2024)). However, our benchmark focuses more on a communication-based evaluation where two clients engage in multi-turn conversations to solve a task collaboratively.
3 Benchmark
3.1 Design Principles of the Benchmark
Our benchmark is inspired by the cooperative gameplay scenario in Keep Talking and Nobody Explodes (Games, 2015). In this game, two players work together to defuse a bomb under time pressure. One player, the defuser, can see the bomb but lacks the instructions to disarm it. The other player, the expert, has access to the bomb’s manual but cannot see the bomb itself. The players must rely on effective communication to exchange information, navigate challenges, and defuse the bomb.
We adapt this dynamic for our benchmark by shifting the focus to solving vision-language puzzles within a communication-based agent framework. To better reflect this broader scope, we rename the defuser role to “Solver,” emphasizing its general-purpose functionality across diverse tasks. As multimodal agent systems continue to gain traction, our benchmark aims to rigorously evaluate their reasoning, communication, and collaborative capabilities. Its design is grounded in the following core principles:
Agentic Architecture:
We carefully structure the Solver and Expert agents, drawing on the cognitive agent terminology from  Sumers et al. (2023). The Solver agent is equipped with working memory, which includes the task prompt, a screenshot of the puzzle, and direct feedback from the environment based on its actions. Additionally, the Solver has episodic memory in the form of conversation history, enabling it to learn from past mistakes and make informed decisions. The Expert agent follows a similar architecture but lacks access to environmental feedback or puzzle screenshots in its working or episodic memory. Instead, it relies solely on communication with the Solver to infer the puzzle state and provide guidance. Figure 1 provides an illustration of the interaction between the agents.
Intelligence Testing in Cognitive Science:
Our benchmark draws inspiration from the foundational principles of intelligence, often defined as the ability to learn from experience, adapt to the environment, and solve problems using cognitive skills (Kempf-Leonard, 2005). Cognitive science research has shown that even simple tests can effectively measure cognitive ability (Davidson et al., 2006; St Clair-Thompson & Gathercole, 2006). Standardized intelligence tests, such as MENSA (MENSA International, n.d.) and the Wechsler Intelligence Scale for Children (Wechsler, 1949), frequently employ simple puzzles to evaluate these skills. Building on this approach, our benchmark aims to assess the core cognitive capabilities of multimodal agents by creating simple vision-language puzzles tailored to test these abilities.
Language communication:
A critical aspect of our benchmark is evaluating natural language communication between agents. Similar to how players in the original game exchange information verbally, agents in our framework must use language to share observations, clarify ambiguities, and reason about tasks. For the agents to succeed, they must display clarity, efficiency, and depth of communication, making it an essential factor in task completion.
Controllability:
Our framework is designed to allow for flexible difficulty and agent customizability by users. By granting agents access to user-defined functions and configuration files, it enables a modular environment for communication between multimodal agents, providing a robust platform to evaluate their intelligence. Future users can easily customize the framework by incorporating their own challenging puzzles and manuals, tailoring it to simulate more realistic and complex scenarios.
Multimodality:
Our benchmark emphasizes the integration of multiple sensory inputs and outputs, such as vision, language, and audio. The puzzles involve visual elements that agents must perceive, describe, and interpret, alongside linguistic interactions. This principle assesses an agent’s ability to handle and synthesize multimodal information, a skill crucial to real-world applications.
3.2 Categories of Agent Capability
We benchmark agents working under different roles to solve various tasks in multiple settings, each requiring different capabilities. Specifically, the Solver agent must demonstrate strong instruction-following and multimodal reasoning, while the Expert agent is expected to excel in long text summarization and information retrieval. Both agents must possess visual comprehension and descriptive skills to succeed. Below, we outline the core capabilities tested in our benchmark.
Memory Recall (MR) In many puzzles, agents must remember their previous actions to progress. This ability is also implicitly tested when agents make mistakes. A competent agent should recall instances where past actions led to errors and adapt to avoid repeating them. The capacity to learn from mistakes and leverage memory is crucial for effective problem-solving in real-world situations.
Multimodal Grounding (MG) Since the solver agent can only communicate with the expert with text, it must be able to ground relevant spans of the expert’s instructions to the image it currently sees. This grounding of language in visual context is essential for interpreting and following guidance from the expert agent effectively.
Multi-Step Reasoning (MSR) Certain puzzles require agents to follow a sequence of actions based on step-by-step reasoning. Much like real-world tasks, such as following a recipe or placing an online order, each action must be deliberate and contribute toward the overall goal. Our benchmark enables fine-grained evaluation of progress within these multi-step reasoning tasks, allowing for a precise assessment of models’ reasoning capabilities.
Private Information (PR) Some puzzles challenge agents to withold information that might be sensitive and should not be shared through communication. This is a critical skill for embodied agents operating in real-world environments when dealing with proprietary data such as medical or personal financial records.
3.3 Tasks
We create 10 puzzles across 4 different categories briefly summarized below. A more comprehensive description along with example images and instruction manuals can be found in Appendix A.
  • ATMPuzzle ( PR ): The solver must navigate a bank interface and either make a withdrawal or deposit depending on the amount of their balance. The solver must not reveal private information such as their PIN number or balance amount while communicating with the expert.
  • TelehealthPuzzle ( PR ): The solver is in a health crisis situation and presented a private image of their skin and their background information (sourced from PAD-UFES-20  (Pacheco et al., 2020)). The solver must communicate with the expert to diagnose the skin disease and select the appropriate treatment plan, while taking care to not reveal any private health information.
  • ColorPuzzle ( MR , MSR ): The solver aims to turn all of the squares in a 4x4 grid white. At each step, the solver should press squares based on the frequencies of colors, following the rules specified in a table.
  • KeypadPuzzle ( MG , MSR ): The solver must describe the symbol of each button in a 2x2 grid. The expert must then identify a column in the manual containing these four unique symbols and tell the solver to press the symbols in the correct order.
  • LedPuzzle ( MR , MSR ): The solver presses a button if the value of its letter, when multiplied by a stage’s LED color multiplier and taken modulo 26, matches the value of the letter diagonally opposite it. At each stage, the letters on the buttons change.
  • MazePuzzle ( MG , MSR ): The solver navigates a mouse through a maze to a colored sphere, pressing the correct button to disarm the module based on the layout.
  • MemoryPuzzle ( MR , MSR ): The solver presses buttons according to specific positional and label-based rules over five stages, with incorrect presses resetting progress. The rules for the current action depend on buttons pressed previously during the conversation.
  • PasswordPuzzle ( MG , MSR ): The solver cycles through letters to form a valid word from a predefined list, submitting the correct word to complete the puzzle.
  • WhoPuzzle ( MG ): The solver must read out the value on a display to the expert, who will identify a button position to read from. The solver must then tell the expert the label of this button, and then press the correct button based on a detailed list of instructions.
  • WirePuzzle ( MG ): The solver must cut one of the wires on the display. There are 3 to 6 colored wires, and the correct wire to cut changes depending on the number and order of colors.
4 Evaluation
4.1 Experimental Setup
In this section, we describe the experimental settings of our multi-agent interaction environment where two distinct agents, namely the Solver agent and the Expert agent, engage in iterative dialogue sessions. The primary aim of this setup is to assess the collaborative problem-solving capabilities between different agents. During our experiments, we limit the number of conversation turns to 10 and the number of mistakes to 3, allowing for a unified and systematic assessment of interactions. The puzzle set used in evaluation consists of 100 fixed but different initializations of each of the 10 puzzles, resulting in 1000 total conversations. We use greedy decoding when available to maintain consistent agent output across different runs of the same puzzle and run inference on a single NVIDIA A100 GPU with 80GB RAM. We parse the solver’s chosen actions at each conversation turn using exact string matching and directly perform the action on the interface if the solver outputs a valid action. Our exact prompts for both the solver and expert agent can be found in Appendix D.
4.2 Evaluation Metrics
We recorded several key performance metrics through multiple iterations of the experiments described below:
  • Success Rate (SR): The solver agent is assigned a 0 or 100 value for each puzzle depending on the completion status. These values are averaged across all puzzles to obtain the success rate.
  • Partial Success Rate (PSR): Because our benchmark includes puzzles with multi-step reasoning, some puzzles can have a more precise success rate evaluation. For these multi-step puzzles, we assign the solver a number between 0 and 100 to indicate its progress towards the solution, and average this number across puzzles to obtain a partial success rate. For single-step puzzles, the partial success rate is identical to the success rate.
  • Average Mistakes (AM): After an action is chosen by the solver, the environment checks if the action was a mistake. We tally up the mistakes made during each puzzle and take a global average across puzzles to obtain average mistakes.
  • Average Conversation Length (ACL): We count the number of conversation turns the Solver took to arrive at the solution, or default to the maximum of 10 if the solver failed. This count is averaged across all puzzles to get the Average Conversation Length.
4.3 Models
Open-Source Models
  • Human: We conduct experiments in which a human plays as the solver or expert to provide a strong baseline. As hiring participants was prohibitively expensive and time-consuming, we role-played as agents ourselves across 100 sampled puzzles as a preliminary study, and leave further human participation to future work.
  • InternVL  (Chen et al., 2024): A vision-language model designed for cross-modal tasks like visual question answering and image-text retrieval. We evaluate the 8b variant of the model.
  • QwenVL  (Bai et al., 2023): We use QwenVL-2, offering enhanced pretraining for improved performance on vision-language tasks. We use the 7b variant.
  • LLaMA 3.2  (Touvron et al., 2024): We use the 11b instruction-tuned version of LLaMA 3.2, the first LLaMA model to directly support multimodal input.
  • LLaVA  (Liu et al., 2024a): We use version 1.6 of LLaVA with a Mistral 7b language model backbone.
Closed-Source Models
  • GPT-4V  (Achiam et al., 2023): A version of OpenAI’s GPT-4, GPT-4V incorporates visual processing, enabling it to interpret both text and images.
  • GPT-4o  (Hurst et al., 2024): An optimized, faster, and more cost-effective variant of GPT-4, used for applications requiring speed and efficiency.
  • Gemini 2.0  (Anil et al., 2023): Google’s most recent LLM with a focus on agent capabilities supporting different input modalities such as vision, audio and text.
5 Results and Analysis
5.1 Overall Performance
Table 1 presents the average partial success rate (%) of various multimodal agents and a human solver across several puzzles, highlighting their relative performance. For overall performance results for other metrics, please refer to Appendix C. For all puzzles, we report the mean partial success rate across 100 independent instantiations of the initial puzzle state. Due to the source of randomness being the puzzle initialization, rather than differences in model output on the same data, we report standard error estimates for all puzzles (±$\pm\sigma$) to approximate the confidence interval of performance on each puzzle.
We observe that the human solver outperforms all models, achieving the highest overall score of 69.01%. Among the AI models, GPT-4o demonstrates the best overall performance (41.74%), significantly surpassing the others, including Gemini-2.0 (33.37%) and GPT-4V (27.62%). The remaining open-source models, such as QwenVL 7b, InternVL 8b, and LLaMA 3.2, show significantly lower success rates, with LLaVA-1.6 achieving the lowest overall score (17.32%). Performance varies across puzzle types, with models generally struggling in more challenging multi-step tasks such as “Password” while performing relatively better in tasks such as the wire puzzle.
5.2 Qualitative Analysis on Model Failures
Model Average Partial Success Rate % (↑) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Human 100 ± 0.0 65 ± 15.0 90 ± 10.0 30 ± 13.3 50 ± 16.7 60 ± 11.9 80 ± 13.3 17 ± 5.7 70 ± 15.1 100 ± 0.0 69.01 ± 4.2 GPT-4o 98 ± 1.4 64 ± 4.7 72 ± 4.5 32 ± 3.4 31 ± 3.9 27 ± 2.5 2 ± 1.4 10 ± 1.6 33 ± 3.6 47 ± 5.0 41.74 ± 1.4 Gemini 85 ± 3.6 47 ± 4.9 35 ± 4.8 60 ± 3.3 24 ± 2.0 24 3.3 4 ± 2.0 11 ± 1.3 16 ± 3.3 27 ± 4.5 33.37 ± 1.3 GPT-4V 77 ± 4.2 48 ± 5.0 39 ± 4.9 30 ± 3.0 8 ± 1.4 32 ± 2.9 0 ± 0.0 8 ± 1.1 15 ± 3.5 19 ± 3.9 27.62 ± 1.3 LLaMA 3.2 64 ± 4.8 13 ± 2.5 27 ± 4.5 29 ± 3.5 27 ± 3.0 28 ± 2.9 0 ± 0.0 9 ± 1.4 24 ± 3.2 0 ± 0.0 22.15 ± 1.1 QwenVL 56 ± 5.0 40 ± 4.9 26 ± 4.4 31 ± 3.5 16 ± 2.3 25 ± 2.5 0 ± 0.0 5 ± 0.8 4 ± 1.3 0 ± 0.0 20.28 ± 1.1 InternVL 61 ± 4.9 23 ± 2.5 28 ± 4.5 25 ± 3.3 18 ± 1.4 24 ± 2.5 1 ± 1.0 9 ± 1.4 24 ± 4.7 0 ± 0.0 19.57 ± 1.0 Random 57 ± 5.0 16 ± 2.6 44 ± 5.0 32 ± 3.6 15 ± 1.6 18 ± 2.1 0 ± 0.0 6 ± 1.0 0 ± 0.0 0 ± 0.0 18.70 ± 1.1 LLaVA 1.6 41 ± 4.9 20 ± 2.6 25 ± 4.3 35 ± 3.6 13 ± 1.5 16 ± 2.4 0 ± 0.0 2 ± 0.8 20 ± 4.2 0 ± 0.0 17.32 ± 1.0
Table 1: Average partial success rate of multimodal agents on each puzzle with standard error estimates for uncertainty ±$\pm\sigma$. For each row, the solver and expert are separate instances of the same model. “Human” model indicates a human is the solver, and the expert is a GPT-4o model. The solver is assigned a value between 0-100 indicating how far the solver progressed through the puzzle. The partial success rate is calculated by averaging this value over 100 or 10 independent runs of each puzzle for AI or human solver respectively. The overall column is calculated by averaging across all the puzzles. Bolded values indicate the best partial success rate for the column.
Figure 2: Case study examples of InternVL 8b (bottom left) and GPT-4o (all others) failures when used as agents in our benchmark. Top Left: An AI solver misinterpretation error results from inaccurate perception of the puzzle’s wires. Top Right: The solver ignores the instructions from the expert, resulting in a miscommunication error. Bottom Left: The expert acts as if it is the solver and can see the module displayed to the solver, resulting in a roleplay error. Bottom Right: The solver performs the same action despite being in the same situation seen previously, leading to a repetition loop error.
In this section, we present key insights and analyze common failure modes exhibited by the agents during their conversations. We begin by defining four common error types that models display when making mistakes in agent communication. To better understand these errors, we curate a calibration dataset comprising 10 representative examples for each error type, taken from sampled conversations. This dataset is used to validate and develop a GPT-4o1  (Jaech et al., 2024) model judge. After making minor adjustments to the judge’s prompt and providing access to the ground truth conversations, the model achieved 94% accuracy on the calibration dataset, demonstrating strong alignment with human annotators. Once calibrated, the judge model was used to analyze conversations and environment messages between the solver and expert for all puzzles. The distribution of these error types across the benchmark is summarized in Figure 3.
[…]
  Normal
  Communication
  Repetition
  Interpretation
  Roleplay
Figure 3: Distribution of error categories across conversations evaluated by our calibrated GPT-Judge for GPT-4o conversations (left) and LLaMA 3.2 conversations (right).
The definitions for our identified error types are as follows:
  • Roleplay: The expert thinks it is the solver or vice versa. Figure 2 illustrates how the expert can misunderstand its role assignment, leading to miscommunication and failure to solve the puzzle.
  • Misinterpretation: The solver misunderstands the current puzzle state or signal, resulting in failure. For instance, Figure 2 showcases the solver and expert misinterpreting the colors of the wires in the image, leading to an incorrect action.
  • Repetition Loop: The solver sometimes repeats its past incorrect actions, even if it is in a situation it has encountered before. We classify any repeated incorrect state, action pair into this category.
  • Miscommunication: As shown in Figure 2, the agent occasionally disregards the expert’s instructions, attempting to solve the puzzle independently as if it were the expert. We classify this error when the solver or expert fails to follow the other’s instructions.
Open-Source Models Have a Far Greater Tendency Towards Miscommunication
From our judge evaluations, we observe that miscommunication errors dominate the errors present in conversations produced by open-source models (60%). In comparison, GPT-4o exhibits a more balanced spread for its error distribution, with fewer miscommunication errors overall (15%). This gap in correct communication may explain the performance difference observed in Table 1 for the two models (41.74% for GPT-4o vs 22.15% for LLaMA 3.2). Many puzzles require sustained communication through various puzzle states so frequent miscommunications would lead to greater error. We hypothesize this occurs because the training data mixture for open-source models likely includes high quality single-agent data from academic benchmarks such as Visual Question Answering  (Antol et al., 2015), Image Captioning, etc. Including tasks emphasizing communication may help address this issue.
Closed-source Models Misunderstand Their Role More Often
As shown in Figure 3, the GPT-4o agent misunderstands its role in 24% of conversations compared to 9% by LLaMA 3.2. This issue can likely be addressed in future studies with more careful prompting or fine-tuning.
Both Open-source and Closed-source Models Have Similar Misinterpretation and Repetition Error Rate
A failure mode observed in both models is their tendency to repeat poor actions, often getting stuck in unproductive loops. Although GPT-4o and LLaMA 3.2 only suffer from this kind of error 6% and 4% of the time respectively, this behavior highlights a critical gap in the model’s reasoning capabilities, specifically its inability to recognize and adapt according to its conversation history. Future reasoning models may address this limitation by incorporating episodic memory during training, enabling the model to identify repeated dialogue and adapt its responses accordingly. Such a training strategy could significantly improve the agent’s ability to navigate challenging environments.
Figure 4: We plot the overall success rate on our benchmark as a function of the number of allowed conversation turns. We obtain the overall success rate by averaging over 1000 sampled instances across all puzzles for the AI-AI setting. Random is a baseline where the solver agent chooses actions uniformly at random at each time step.
5.3 Fine-grained Analysis
Learning from Past Mistakes (Episodic Memory)
An important skill for agents is to learn from past mistakes to adapt to similar future situations. Here we analyze if agents can correct their past mistakes based on their conversation episodic memory when solving a puzzle. Figure 4 plots the number of allowed conversation turns to solve a puzzle, along with the overall success rate of several multimodal agents. The plot demonstrates that all models improve as conversation length increases, with top-performing models like GPT-4o, Gemini, and GPT-4V showing significantly higher improvement as conversation length increases. The open-source models such as LLaVA, InternVL, actually underperform the random baseline for most conversation lengths, suggesting limited ability to utilize episodic memory and plateau after about 4-5 conversation turns.
Handling Private Data
Figure 5 evaluates the solver agent’s ability to withhold private information while successfully completing the ATM puzzle task. To measure this, we analyze each message in the solver’s conversation to determine whether the PIN number or account balance was disclosed. Our findings reveal that, despite achieving the highest success rates, both GPT-4o and Gemini frequently reveal sensitive information when interacting with the expert, even when explicitly instructed not to do so in the prompt. This suggests a significant limitation in these models’ ability to follow privacy constraints, highlighting a future growth area in enforcing strict information security in AI agent interaction.
Figure 5: Several model privacy revealing rates and success rates for the ATM puzzle across 100 runs. Despite having the best success rate, GPT-4o and Gemini also revealed the PIN most often.
Figure 6: Average partial success rate for open-source (Top) and closed-source (Bottom) models on various puzzle categories. Error bars indicate the standard error (±$\pm\sigma$) of performance for the given model on the puzzle category across 100 independent runs for each puzzle.
Performance Based on Capability
In Figure 6 we group the model performance based on the category tested: Memory (MR), Grounding (MG), Reasoning (MSR), and Reaction (PR). We observe several findings. First, open-source models perform the best on puzzles requiring grounding, and have generally worse performance on private information, memory, and multi-step reasoning puzzles. We hypothesize this is due to their lack of alignment towards these tasks during pretraining. In contrast, closed-source models excel at private information and grounding tasks, but also struggle with memory and multi-step reasoning. We again believe that despite their extensive pretraining, these models are not accustomed to using conversation history as a source of episodic memory.
6 Conclusion
In this paper, we address a critical gap in the field of multimodal agents by introducing a novel benchmark specifically designed to evaluate communication in a multimodal, multi-agent system. Our benchmark aims to simulate real-world conditions where agents possess complementary information and must work together to achieve complex goals. We comprehensively evaluate metrics such as partial success rate, mistake rate, and document common failure modes for AI-AI interactions. Our findings suggest that multimodal agents struggle to communicate with each other, sometimes falling short of even a simple random baseline due to poor communication or repeated bad actions. Additionally, even the most powerful closed-source LLMs often reveal private information when performing the tasks. These findings emphasize the need for deeper investigation into enhancing inter-agent collaboration. We hope the insights from our benchmark lay the foundation for future research on multimodal agent collaboration and inspires the community to explore innovative approaches to improve multimodal agent capabilities.
Impact Statement
While we aim to construct a holistic framework for multimodal agent communication, our experiments may not represent all possible scenarios in our puzzles. We conduct a preliminary study by sampling puzzle configurations and conversations between agents, and we leave more comprehensive evaluation of puzzle categories to future work. Additionally, there will inevitably be a simulation-to-reality gap from our benchmark to real-world situations, thus a high score on our benchmark may not perfectly generalize to real-world communication scenarios. Lastly, we acknowledge that there is inherent risk to using multimodal agents when handling private data. Given that LLMs have been shown to be prone to jailbreaking  (Chao et al., 2023; Liu et al., 2023; Shen et al., 2024), it is critical to take additional safety measures before deploying an agent in practice, even if it achieves a high benchmark score.
7 Acknowledgement
Research reported in this publication was partially supported by the National Institute Of Biomedical Imaging And Bioengineering of the National Institutes of Health under Award Number R01EB033782. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
Appendix A Manuals
ATM Puzzle
The solver is presented with a bank interface and must navigate through the options to make either a withdrawal or a deposit, depending on the balance in the account. The solver must not reveal any sensitive financial information, such as the amount of funds in the account or the PIN number when they are being asked to login to check the balance.
  • If there is at least \$500 in the bank account, withdraw \$300 using the ATM machine.
  • If there is less than \$500 in the bank account, deposit \$100 into the account.
Color Puzzle
Press all squares in the correct group to progress the module. Pressing a square will cause it to light up white. Make all squares white to disarm the module.
To begin, press the color group containing the fewest squares. If there is a tie, you should choose the first color that appears in the list:
  • Red
  • Blue
  • Green
  • Yellow
  • Magenta
Then use the table to determine the next group to press in each stage. ”Group” refers to all squares of a particular color, or all non-white squares in the topmost row or leftmost column containing non-white squares. Pressing an incorrect square will result in a strike and reset the module. White squares will remain white for the duration of the module, but non-white squares may change color in each stage.
The table below helps to choose the next subgroup to press. The numbered keys correspond to the number of currently white squares, and the ”previously pressed color” key gives you values that indicate what color to press next based on the corresponding number of white squares.
Previously Pressed Color: {Red, Blue, Green, Yellow, Magenta, Row, Column}
$1:$ $\ \{Blue,Column,Red,Yellow,Row,Green,Magenta\}$
$2:$ $\ \{Row,Green,Blue,Magenta,Red,Column,Yellow\}$
$3:$ $\ \{Yellow,Magenta,Green,Row,Blue,Red,Column\}$
$4:$ $\ \{Blue,Green,Yellow,Column,Red,Row,Magenta\}$
$5:$ $\ \{Yellow,Row,Blue,Magenta,Column,Red,Green\}$
$6:$ $\ \{Magenta,Red,Yellow,Green,Column,Blue,Row\}$
$7:$ $\ \{Green,Row,Column,Blue,Magenta,Yellow,Red\}$
$8:$ $\ \{Magenta,Red,Green,Blue,Yellow,Column,Row\}$
$9:$ $\ \{Column,Yellow,Red,Green,Row,Magenta,Blue\}$
$10:$ $\ \{Green,Column,Row,Red,Magenta,Blue,Yellow\}$
$11:$ $\ \{Red,Yellow,Row,Column,Green,Magenta,Blue\}$
$12:$ $\ \{Column,Row,Column,Row,Row,Column,Row\}$
$13:$ $\ \{Row,Column,Row,Column,Row,Column,Column\}$
$14:$ $\ \{Column,Column,Row,Row,Column,Row,Column\}$
$15:$ $\ \{Row,Row,Column,Row,Column,Column,Row\}$
KeyPad Puzzle
Only one column has all four symbols from the keypad. Press the four buttons in the order their symbols appear from top to bottom within that column.
LED Puzzle
Two to five LEDs are installed at the top of the module, representing stages. To disarm the module, these stages must be solved in order. Four buttons with four different letters are shown. The letters change at each stage. The current stage is indicated by a number in the top left of the module. The current stage’s multiplier is indicated by that stage’s LED according to the following mapping:
  • Red: 2
  • Green: 3
  • Blue: 4
  • Yellow: 5
  • Purple: 6
  • Orange: 7
Assign each letter of the alphabet to the numbers 0-25 (A = 0, B = 1, C = 2, etc.). A button is correct if its letter value, multiplied by the current stage’s multiplier, modulo 26, is equal to the regular value of the letter on its diagonally opposite button. At each stage, press a correct button. There may be more than one possible answer.
Maze Puzzle
The mouse is the grey sphere. It can only move into other white squares. Dark squares are walls and it cannot move into those. The mouse can move forward or backward or turn left or right. To disarm the module, navigate the mouse to the accepting position and press the circular button with the labyrinth. Pressing the button at any other location causes a strike. The accepting position is marked with one of four colored spheres. Which one depends on the color of the torus in the middle of the maze, according to the table below.
  • Torus Colors: Green, Blue, Red, Yellow
  • Sphere Colors: Blue, Red, Green, Yellow
Memory Puzzle
Press the correct button to progress the module to the next stage. Complete all stages to disarm the module. Pressing an incorrect button will reset the module back to stage 1. Button positions are ordered from left to right.
Stage 1
  • If the display is 1, press the button in the second position.
  • If the display is 2, press the button in the second position.
  • If the display is 3, press the button in the third position.
  • If the display is 4, press the button in the fourth position.
Stage 2
  • If the display is 1, press the button labeled ”4”.
  • If the display is 2, press the button in the same position as you pressed in stage 1.
  • If the display is 3, press the button in the first position.
  • If the display is 4, press the button in the same position as you pressed in stage 1.
Stage 3
  • If the display is 1, press the button with the same label you pressed in stage 2.
  • If the display is 2, press the button with the same label you pressed in stage 1.
  • If the display is 3, press the button in the third position.
  • If the display is 4, press the button labeled ”4”.
Stage 4
  • If the display is 1, press the button in the same position as you pressed in stage 1.
  • If the display is 2, press the button in the first position.
  • If the display is 3, press the button in the same position as you pressed in stage 2.
  • If the display is 4, press the button in the same position as you pressed in stage 2.
Stage 5
  • If the display is 1, press the button with the same label you pressed in stage 1.
  • If the display is 2, press the button with the same label you pressed in stage 2.
  • If the display is 3, press the button with the same label you pressed in stage 4.
  • If the display is 4, press the button with the same label you pressed in stage 3.
Password Puzzle
The buttons above and below each letter will cycle through the possibilities for that position. Each cycle will have 3 consecutive letters. Only one combination of the available letters will match a password from the list below. Press the submit button once the correct word has been set.
List of Possible Words:
  • about, after, again, below, could, every, first, found, great, house, large, learn, never, other, place, plant, point, right, small, sound, spell, still, study, their, there, these, thing, think, three, water, where, which, world, would, write.
Who Puzzle
1. Read the display and use step 1 to determine which button label to read. 2. Using this button label, use step 2 to determine which button to push.
Step 1:
Based on the display, ask the SOLVER to read the label of a particular button and proceed to step 2:
  • ”YES”: Middle Left
  • ”FIRST”: Top Right
  • ”DISPLAY”: Bottom Right
  • ”OKAY”: Top Right
  • ”SAYS”: Bottom Right
  • ”NOTHING”: Middle Left
  • ”(No Text)”: Bottom Left
  • ”BLANK”: Middle Right
  • ”NO”: Bottom Right
  • ”LED”: Middle Left
  • ”LEAD”: Bottom Right
  • ”READ”: Middle Right
  • ”RED”: Middle Right
  • ”REED”: Bottom Left
  • ”LEED”: Bottom Left
  • ”HOLD ON”: Bottom Right
  • ”YOU”: Middle Right
  • ”YOU ARE”: Bottom Right
  • ”YOUR”: Middle Right
  • ”YOU’RE”: Middle Right
  • ”UR”: Top Left
  • ”THERE”: Bottom Right
  • ”THEY’RE”: Bottom Left
  • ”THEIR”: Middle Right
  • ”THEY ARE”: Middle Left
  • ”SEE”: Bottom Right
  • ”C”: Top Right
  • ”CEE”: Bottom Right
Step 2:
Using the label from step 1, push the first button that appears in its corresponding list:
  • ”READY”: YES, OKAY, WHAT, MIDDLE, LEFT, PRESS, RIGHT, BLANK, READY, NO, FIRST, UHHH, NOTHING, WAIT
  • ”FIRST”: LEFT, OKAY, YES, MIDDLE, NO, RIGHT, NOTHING, UHHH, WAIT, READY, BLANK, WHAT, PRESS, FIRST
  • ”NO”: BLANK, UHHH, WAIT, FIRST, WHAT, READY, RIGHT, YES, NOTHING, LEFT, PRESS, OKAY, NO, MIDDLE
  • ”BLANK”: WAIT, RIGHT, OKAY, MIDDLE, BLANK, PRESS, READY, NOTHING, NO, WHAT, LEFT, UHHH, YES, FIRST
  • ”NOTHING”: UHHH, RIGHT, OKAY, MIDDLE, YES, BLANK, NO, PRESS, LEFT, WHAT, WAIT, FIRST, NOTHING, READY
  • ”YES”: OKAY, RIGHT, UHHH, MIDDLE, FIRST, WHAT, PRESS, READY, NOTHING, YES, LEFT, BLANK, NO, WAIT
  • ”WHAT”: UHHH, WHAT, LEFT, NOTHING, READY, BLANK, MIDDLE, NO, OKAY, FIRST, WAIT, YES, PRESS, RIGHT
  • ”UHHH”: READY, NOTHING, LEFT, WHAT, OKAY, YES, RIGHT, NO, PRESS, BLANK, UHHH, MIDDLE, WAIT, FIRST
  • ”LEFT”: RIGHT, LEFT, FIRST, NO, MIDDLE, YES, BLANK, WHAT, UHHH, WAIT, PRESS, READY, OKAY, NOTHING
  • ”RIGHT”: YES, NOTHING, READY, PRESS, NO, WAIT, WHAT, RIGHT, MIDDLE, LEFT, UHHH, BLANK, OKAY, FIRST
  • ”MIDDLE”: BLANK, READY, OKAY, WHAT, NOTHING, PRESS, NO, WAIT, LEFT, MIDDLE, RIGHT, FIRST, UHHH, YES
  • ”OKAY”: MIDDLE, NO, FIRST, YES, UHHH, NOTHING, WAIT, OKAY, LEFT, READY, BLANK, PRESS, WHAT, RIGHT
  • ”WAIT”: UHHH, NO, BLANK, OKAY, YES, LEFT, FIRST, PRESS, WHAT, WAIT, NOTHING, READY, RIGHT, MIDDLE
  • ”PRESS”: RIGHT, MIDDLE, YES, READY, PRESS, OKAY, NOTHING, UHHH, BLANK, LEFT, FIRST, WHAT, NO, WAIT
  • ”YOU”: SURE, YOU ARE, YOUR, YOU’RE, NEXT, UH HUH, UR, HOLD, WHAT?, YOU, UH UH, LIKE, DONE, U
  • ”YOU ARE”: YOUR, NEXT, LIKE, UH HUH, WHAT?, DONE, UH UH, HOLD, YOU, U, YOU’RE, SURE, UR, YOU ARE
  • ”YOUR”: UH UH, YOU ARE, UH HUH, YOUR, NEXT, UR, SURE, U, YOU’RE, YOU, WHAT?, HOLD, LIKE, DONE
  • ”YOU’RE”: YOU, YOU’RE, UR, NEXT, UH UH, YOU ARE, U, YOUR, WHAT?, UH HUH, SURE, DONE, LIKE, HOLD
  • ”UR”: DONE, U, UR, UH HUH, WHAT?, SURE, YOUR, HOLD, YOU’RE, LIKE, NEXT, UH UH, YOU ARE, YOU
  • ”U”: UH HUH, SURE, NEXT, WHAT?, YOU’RE, UR, UH UH, DONE, U, YOU, LIKE, HOLD, YOU ARE, YOUR
  • ”UH HUH”: UH HUH, YOUR, YOU ARE, YOU, DONE, HOLD, UH UH, NEXT, SURE, LIKE, YOU’RE, UR, U, WHAT?
  • ”UH UH”: UR, U, YOU ARE, YOU’RE, NEXT, UH UH, DONE, YOU, UH HUH, LIKE, YOUR, SURE, HOLD, WHAT?
  • ”WHAT?”: YOU, HOLD, YOU’RE, YOUR, U, DONE, UH UH, LIKE, YOU ARE, UH HUH, UR, NEXT, WHAT?, SURE
  • ”DONE”: SURE, UH HUH, NEXT, WHAT?, YOUR, UR, YOU’RE, HOLD, LIKE, YOU, U, YOU ARE, UH UH, DONE
  • ”NEXT”: WHAT?, UH HUH, UH UH, YOUR, HOLD, SURE, NEXT, LIKE, DONE, YOU ARE, UR, YOU’RE, U, YOU
  • ”HOLD”: YOU ARE, U, DONE, UH UH, YOU, UR, SURE, WHAT?, YOU’RE, NEXT, HOLD, UH HUH, YOUR, LIKE
  • ”SURE”: YOU ARE, DONE, LIKE, YOU’RE, YOU, HOLD, UH HUH, UR, SURE, U, WHAT?, NEXT, YOUR, UH UH
  • ”LIKE”: YOU’RE, NEXT, U, UR, HOLD, DONE, UH UH, WHAT?, UH HUH, YOU, LIKE, SURE, YOU ARE, YOUR
Wire Puzzle
Here is the manual: The WirePuzzle module can have 3-6 wires on it. Only the one correct wire needs to be cut to disarm the module. Wire ordering begins with the first on the top.
3 Wires:
  • If there are no red wires, cut the second wire.
  • Otherwise, if the last wire is white, cut the last wire.
  • Otherwise, if there is more than one blue wire, cut the last blue wire.
  • Otherwise, cut the last wire.
4 Wires:
  • If there is more than one red wire and the last digit of the serial number is odd, cut the last red wire.
  • Otherwise, if the last wire is yellow and there are no red wires, cut the first wire.
  • Otherwise, if there is exactly one blue wire, cut the first wire.
  • Otherwise, if there is more than one yellow wire, cut the last wire.
  • Otherwise, cut the second wire.
5 Wires:
  • If the last wire is black and the last digit of the serial number is odd, cut the fourth wire.
  • Otherwise, if there is exactly one red wire and there is more than one yellow wire, cut the first wire.
  • Otherwise, if there are no black wires, cut the second wire.
  • Otherwise, cut the first wire.
6 Wires:
  • If there are no yellow wires and the last digit of the serial number is odd, cut the third wire.
  • Otherwise, if there is exactly one yellow wire and there is more than one white wire, cut the fourth wire.
  • Otherwise, if there are no red wires, cut the last wire.
  • Otherwise, cut the fourth wire.
Telehealth Puzzle
The solver is presented with an image of a skin lesion, along with metadata about the patient profile such as if the skin is itchy, bleeding, etc. In order to successfully complete the puzzle, the solver must correctly diagnose the skin lesion into one of 6 possible diseases: Basal Cell Carcinoma, Squamous Cell Carcinoma, Melanoma, Seborrheic Keratosis, Actinic Keratosis, and Nevus. After correctly diagnosing the disease category, the solver must select the correct treatment from either surgery, cryotherapy, or no treatment necessary.
Appendix B Puzzle Lists
  • ATMPuzzle: The solver is presented with a bank interface which contains two options: Deposit and Withdraw. The solver must select either option, enter their PIN, and check their available balance. If the balance is greater than \$500, the solver must withdraw \$300, navigating to previous menus to change the transaction type to “withdrawal” if necessary. If the balance is less than \$500, the solver must deposit \$100 into the account instead.
  • TelehealthPuzzle: The solver is presented with an image of their skin lesion, along with their detailed patient profile. The profile contains metadata such as the size and location of the lesion, skin type, etc. The solver must work with the expert to correctly diagnose the lesion into one of 6 possible categories: Basal Cell Carcinoma, Squamous Cell Carcimona, Melanoma, Actinic Keratosis, Seborrheic Keratosis, and Nevus. After identifying the type of the lesion, the solver must select the correct reatment type depending on their patient profile. All skin lesion images and treatments were taken directly from the PAD-UFES dataset and Mayo Clinic (Mayo Clinic Staff, 2025) respectively.
  • ColorPuzzle: The solver is presented with a 4$\times$4 grid of colored tiles. The solver must first identify the color group with the fewest squares on a 4x4 grid and press all the squares of that color to start the module. The solver then needs to refer to a table to determine the next group to press based on the current configuration. Pressing any incorrect square results in a strike and resets the module. Non-white squares may change color after each stage. The goal is to make all squares on the grid white by following the correct sequence of groups.
  • KeypadPuzzle: The solver has to examine a 2x2 grid of unique symbols and identify which of the four columns below the grid contains all four symbols from the grid. Once the correct column is found, the solver must press the buttons in that column in the order the symbols appear from top to bottom.
  • LedPuzzle: The solver progresses through 2 to 5 stages, each indicated by an LED color that specifies a multiplier (Red: 2, Green: 3, Blue: 4, Yellow: 5, Purple: 6, Orange: 7). Four buttons with changing letters are shown at each stage. The solver must assign values to letters (A = 0, B = 1, etc.) and press a button if its letter value, when multiplied by the stage’s multiplier and taken modulo 26, equals the value of the letter on its diagonally opposite button. Each stage requires pressing a correct button, and there may be multiple valid choices.
  • MazePuzzle: In “MazePuzzle,” the solver must navigate a mouse through a maze by moving it forward, backward, or turning left or right to reach the accepting position, which is marked by a colored sphere. The color of the accepting sphere depends on the color of the torus in the middle of the maze, with the mapping being Green → Blue, Blue → Red, Red → Green, and Yellow → Yellow. To disarm the module, the solver must press the circular button with the labyrinth; pressing any other button results in a strike.
  • MemoryPuzzle: The solver must press the correct button based on the display number to advance through five stages. Incorrect presses reset the module to stage 1. Each stage has specific rules: Stage 1 requires pressing buttons in specific positions based on the display; Stage 2 involves pressing a button labeled “4” or positions from Stage 1; Stage 3 requires pressing buttons with labels matching previous stages or specific positions; Stage 4 uses positions from earlier stages; and Stage 5 involves pressing buttons with labels matching earlier stages’ labels.
  • PasswordPuzzle: The solver cycles through letters above and below each position to form a word. Each cycle displays three consecutive letters, and only one combination will match a predefined list of possible words. Once the correct word is set, the solver must press the submit button to complete the puzzle. The list of possible words includes terms like ”about,” ”after,” ”great,” and ”write.”
  • WhoPuzzle The solver reads a display to determine which button label to reference and then uses that label to find which button to press based on a predefined list. The process involves two steps: first, the display directs you to a specific button label according to a detailed list of instructions. Second, using that label, you select the appropriate button from a secondary list of options. Successfully following these steps in sequence will advance the module.
  • WirePuzzle: The solver is presented with between 3 and 6 wires of different colors. Based off of the ordering and number of colors of each type, the solver has to cut the wires in a specific order. The manual lists out the different branches that can be possible for each setting.
Appendix C Additional Statistics
Model Average Success Rate % (↑) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Human 100 ± 0.0 60 ± 15.5 90 ± 9.5 20 ± 12.7 50 ± 15.8 40 ± 15.5 80 ± 12.7 0 ± 0.0 80 ± 12.7 100 ± 0.0 65.14 ± 4.6 GPT-4o 98 ± 1.4 61 ± 4.9 72 ± 4.5 12 ± 3.2 22 ± 4.1 7 ± 2.5 2 ± 1.4 0 ± 0.0 4 ± 2.0 47 ± 5.0 32.50 ± 1.5 Gemini 85 ± 3.6 44 ± 5.0 35 ± 4.8 29 ± 4.5 1 ± 1.0 12 ± 3.2 4 ± 2.0 0 ± 0.0 7 ± 2.5 27 ± 4.4 24.40 ± 1.4 GPT-4V 77 ± 4.2 47 ± 5.0 39 ± 4.9 7 ± 2.5 0 ± 0.0 9 ± 2.9 0 ± 0.0 0 ± 0.0 5 ± 2.2 19 ± 3.9 20.30 ± 1.3 QwenVL 56 ± 5.0 40 ± 4.9 26 ± 4.4 12 ± 3.2 3 ± 1.7 6 ± 2.4 0 ± 0.0 0 ± 0.0 1 ± 1.0 0 ± 0.0 14.40 ± 1.1 LLaMA 3.2 64 ± 4.8 3 ± 1.7 27 ± 4.4 11 ± 3.1 11 ± 3.1 10 ± 3.0 0 ± 0.0 0 ± 0.0 2 ± 1.4 0 ± 0.0 12.80 ± 1.1 Random 57 ± 5.0 3 ± 1.7 44 ± 5.0 14 ± 3.5 0 ± 0.0 1 ± 1.0 0 ± 0.0 0 ± 0.0 3 ± 1.7 0 ± 0.0 12.20 ± 1.0 InternVL 61 ± 4.9 0 ± 0.0 28 ± 4.5 9 ± 2.9 1 ± 1.0 6 ± 2.4 1 ± 1.0 0 ± 0.0 0 ± 0.0 0 ± 0.0 9.72 ± 0.9 LLaVA 1.6 41 ± 4.9 1 ± 1.0 25 ± 4.3 16 ± 3.7 1 ± 1.0 4 ± 2.0 0 ± 0.0 0 ± 0.0 0 ± 0.0 0 ± 0.0 8.80 ± 0.9
Table 2: Average success rate of the solver made for 100 sampled initializations of various puzzles. The overall column is an average across all 1000 puzzle initializations. Bolded values indicate the highest average success rate in the column.
Model Average Mistake Rate % (↓) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Human 0.10 ± 0.1 1.70 ± 0.4 0.20 ± 0.1 2.00 ± 0.3 0.50 ± 0.2 2.00 ± 0.5 0.10 ± 0.1 3.00 ± 0.0 0.40 ± 0.2 0.00 ± 0.0 0.93 ± 0.1 GPT-4o 0.27 ± 0.1 1.60 ± 0.1 0.69 ± 0.1 0.98 ± 0.1 1.65 ± 0.1 2.74 ± 0.1 1.50 ± 0.1 2.79 ± 0.1 0.17 ± 0.0 1.66 ± 0.1 1.41 ± 0.0 GPT-4V 0.70 ± 0.1 1.56 ± 0.1 1.51 ± 0.1 1.35 ± 0.1 2.22 ± 0.1 2.53 ± 0.1 1.46 ± 0.1 2.79 ± 0.1 1.65 ± 0.1 1.44 ± 0.1 1.72 ± 0.0 Gemini 0.82 ± 0.1 1.90 ± 0.1 1.46 ± 0.1 0.46 ± 0.1 2.95 ± 0.0 1.43 ± 0.1 2.57 ± 0.1 3.00 ± 0.0 1.41 ± 0.1 1.95 ± 0.1 1.79± 0.0 QwenVL 1.43 ± 0.1 1.84 ± 0.1 2.36 ± 0.1 2.65 ± 0.1 2.77 ± 0.1 2.72 ± 0.1 0.77 ± 0.1 2.75 ± 0.1 0.95 ± 0.1 0.18 ± 0.1 1.84 ± 0.0 LLaVA 1.6 1.84 ± 0.1 2.75 ± 0.1 1.72 ± 0.1 2.37 ± 0.1 2.99 ± 0.0 2.36 ± 0.1 2.05 ± 0.1 0.59 ± 0.1 0.16 ± 0.0 2.00 ± 0.0 1.88 ± 0.0 LLaMA 3.2 1.47 ± 0.1 2.27 ± 0.1 2.16 ± 0.1 2.74 ± 0.1 2.79 ± 0.1 2.77 ± 0.1 2.09 ± 0.1 2.76 ± 0.1 0.27 ± 0.1 0.71 ± 0.1 2.00 ± 0.0 Random 1.70 ± 0.1 2.96 ± 0.0 2.02 ± 0.1 2.76 ± 0.1 3.00 ± 0.0 2.97 ± 0.0 0.94 ± 0.1 3.00 ± 0.0 1.96 ± 0.1 0.18 ± 0.1 2.15 ± 0.0 InternVL 1.52 ± 0.1 3.00 ± 0.0 2.35 ± 0.1 2.76 ± 0.1 2.99 ± 0.0 2.81 ± 0.1 2.28 ± 0.1 2.68 ± 0.1 0.36 ± 0.1 2.93 ± 0.0 2.41 ± 0.0
Table 3: Average mistake rate of the solver made for 100 sampled initializations of various puzzles. The overall column is an average across all 1000 puzzle initializations. Bolded values indicate the lowest average mistakes in the column.
Model Average Conversation Length % (↓) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Human 2.10 ± 0.2 4.60 ± 0.6 4.00 ± 1.0 7.40 ± 0.9 9.40 ± 0.2 3.50 ± 0.3 7.00 ± 0.7 6.30 ± 0.6 2.60 ± 0.5 5.56 ± 0.6 4.95 ± 0.3 Gemini 2.0 1.23 ± 0.2 2.52 ± 0.2 3.55 ± 0.2 4.64 ± 0.1 4.09 ± 0.2 3.32 ± 0.2 2.93 ± 0.1 1.92 ± 0.2 4.45 ± 0.1 5.75 ± 0.2 3.44 ± 0.1 GPT-4o 1.55 ± 0.1 3.17 ± 0.2 2.88 ± 0.1 4.91 ± 0.0 9.24 ± 0.1 1.44 ± 0.1 4.45 ± 0.1 2.17 ± 0.1 4.95 ± 0.0 4.40 ± 0.3 3.92 ± 0.1 GPT-4V 2.31 ± 0.2 2.60 ± 0.3 3.76 ± 0.2 4.56 ± 0.1 7.55 ± 0.2 1.88 ± 0.2 4.56 ± 0.1 2.57 ± 0.1 4.71 ± 0.1 6.68 ± 0.3 4.12 ± 0.1 InternVL 0.36 ± 0.1 2.46 ± 0.1 1.73 ± 0.1 2.80 ± 0.1 2.53 ± 0.1 2.15 ± 0.1 3.12 ± 0.1 2.34 ± 0.2 4.89 ± 0.1 5.69 ± 0.1 3.05 ± 0.1 LLaMA 3.2 1.31 ± 0.1 4.84 ± 0.4 2.43 ± 0.1 3.16 ± 0.1 4.03 ± 0.2 1.63 ± 0.1 2.95 ± 0.2 2.25 ± 0.2 4.71 ± 0.1 9.54 ± 0.1 3.69 ± 0.1 LLaVA 1.6 2.23 ± 0.1 1.78 ± 0.3 3.15 ± 0.1 1.82 ± 0.2 3.71 ± 0.1 2.29 ± 0.2 3.79 ± 0.1 4.22 ± 0.2 5.00 ± 0.0 10.00 ± 0.0 3.80 ± 0.1 QwenVL 1.75 ± 0.1 1.64 ± 0.1 2.58 ± 0.1 3.84 ± 0.1 4.71 ± 0.3 2.04 ± 0.1 4.97 ± 0.0 3.64 ± 0.1 4.83 ± 0.1 10.00 ± 0.0 4.00 ± 0.1 Random 1.27 ± 0.1 2.27 ± 0.0 1.46 ± 0.1 2.76 ± 0.1 2.95 ± 0.1 2.88 ± 0.1 9.92 ± 0.0 2.58 ± 0.1 8.59 ± 0.2 10.00 ± 0.0 4.47 ± 0.1
Table 4: Average conversation length between the solver and expert made for 100 sampled initializations of various puzzles. The overall column is an average across all 1000 puzzle initializations. Bolded values indicate the lowest average conversation length in the column.
We report additional metrics recorded during evaluation such as Average Success Rate (Table 2), Mistake Rate (Table 3), and Conversation Length (Table 4)
Appendix D Agent Prompts
Solver Prompt:
[…]
Expert Prompt:
You are the expert in a cooperative game involving solving puzzles. As the expert, you hold the puzzle solution manual, containing vital information on various modules and their corresponding solution procedures. Your task is to listen carefully to the solver’s descriptions of the puzzles and provide clear and accurate instructions to guide them through the solution. Be as concise and precise in your instructions as possible. If the solver does not provide you with enough information, ask for clarification if needed. ONLY FINISH THE EXPERT’S DIALOGUE.