### v3

# COMMA: A Communicative Multimodal Multi-Agent Benchmark

Abstract
The rapid advances of multimodal agents built on large foundation models have largely overlooked their potential for language-based communication between agents in collaborative tasks. This oversight presents a critical gap in understanding their effectiveness in real-world deployments, particularly when communicating with humans. Existing agentic benchmarks fail to address key aspects of inter-agent communication and collaboration, particularly in scenarios where agents have unequal access to information and must work together to achieve tasks beyond the scope of individual capabilities. To fill this gap, we introduce COMMA: a novel puzzle benchmark designed to evaluate the collaborative performance of multimodal multi-agent systems through language communication. Our benchmark features a variety of multimodal puzzles, providing a comprehensive evaluation across four key categories of agentic capability in a communicative collaboration setting. Our findings reveal surprising weaknesses in state-of-the-art models, including strong proprietary models like GPT-4o and reasoning models like o4-mini. Many chain of thought reasoning models such as R1-Onevision and LLaVA-CoT struggle to outperform even a random baseline in agent-agent collaboration, indicating a potential growth area in their communication abilities.  [^1] Our data and code is available at https://github.com/tossowski/COMMA
1 Introduction
The field of multimodal agents is experiencing rapid growth  (Xu et al., 2024c; Xie et al., 2024; Cao et al., 2024), with research efforts expanding at an unprecedented pace. However, amidst this growth, a critical gap in research has emerged: the lack of focus on collaborative work (Gurcan, 2024; Park et al., 2023; Hong et al., 2024; Liu et al., 2024b) among multiple multimodal agents. Synergistic operation of such agents is a highly promising but largely unexplored domain. Language agents can collaboratively finish complex tasks such as software development  (Qian et al., 2024; Du et al., 2024) or even machine learning research  (Schmidgall et al., 2025) by assuming functional roles such as system designer, function generator, etc. Current research on multimodal agents (Xu et al., 2024c; Xie et al., 2024; Cao et al., 2024) has mainly focused on individual agent capabilities, neglecting the potential for inter-agent collaboration. This limitation is further compounded by existing benchmarks such as TheAgentCompany (Xu et al., 2024a), VisualWebArena (Koh et al., 2024) and MME-RealWorld (Zhang et al., 2024), which do not assess collaborative performance between agents. As a result, our ability to evaluate and improve multi-agent systems remains constrained, hindering progress in this crucial area.
Several critical questions emerge in the context of multimodal agent collaboration. How can different agents effectively communicate multimodal information through language when they have varying levels of access to information? In scenarios where different agents possess diverse task-specific capabilities, how can they collaborate to accomplish objectives beyond the scope of any individual agent? These research settings remain largely uncharted and present significant challenges. Furthermore, the ability of agents to handle incomplete information is of paramount importance, particularly when working with sensitive data  (Li et al., 2024) (i.e. Agent application in healthcare where privacy concerns are critical  (Tang et al., 2024)). Exploration of these questions is crucial for advancing the field of multimodal agent collaboration. By addressing these challenges, we can expand the applicability of multimodal agents in real-world scenarios (Zhang et al., 2024), particularly those involving sensitive or restricted information.
Figure 1: Overview of the interaction between the Solver and Expert agents in our benchmark. Both agents operate with structured input corresponding to working and episodic memory. The Solver receives an image of the puzzle state (working memory) and makes decisions based on the available actions described in the task prompt. The Expert, guided by instruction manuals (working memory), provides advice based on the Solver’s descriptions, such as indicating which buttons to press. The Solver can choose to execute actions by interacting with the environment or communicate with the Expert for further guidance. Their interaction is documented through a dialogue, showcasing the cooperation required to complete the task. Both agents engage in self-reflection by referencing the conversation history, which is continuously updated and included in their input as episodic memory. We include a more comprehensive example in Appendix A
Motivated by these aforementioned issues, we propose a benchmark for evaluating multimodal models in a collaborative agentic setting to address critical gaps in current approaches (see  Figure 1). While recent work has explored scaling test time compute  (Snell et al., 2025) to improve reasoning performance, such methods often rely on the unrealistic assumption that a model has sufficient capability and access to all necessary information to complete the task. In contrast, our benchmark poses a unique challenge: models must dynamically alternate between intensive reasoning based on the information available and effective communication to request additional details when needed. Our evaluation setting also simulates a scenario where an in-house agent with direct access to sensitive data (i.e., the AI solver) collaborates with external expert agents (i.e., the AI expert) to analyze information without compromising privacy. Specifically, this work makes the following contributions:
  • We propose an evaluation benchmark called COMMA, a multimodal agent benchmark which uses puzzles to test language communication between multiple agents (Section 3). To achieve a high score, agents must engage in concise, but meaningful dialogue to arrive at each solution.
  • Using COMMA, we carefully record conversations and performance metrics between state-of-the-art multimodal models such as o4-mini, GPT-4o, Gemini, QwenVL, etc (Section 4).
  • We categorize the agent capabilities tested in our model and common failure modes, providing insight into future research directions for improving inter-agent communication (Section 5).
2 Related Work
Multi-agent Frameworks:
There are many emergent agent collaboration works  (Gurcan, 2024; Park et al., 2023; Hong et al., 2024; Liu et al., 2024b; Ghafarollahi & Buehler, 2024; Li et al., 2023; Wu et al., 2023) among multiple language agents. Multi-agent systems arise mainly in two different scenarios: (1) role-playing different task executors (e.g., software development requiring different roles of agents, such as program manager, software architect, programmer  (Du et al., 2024; Qian et al., 2024; Hong et al., 2024), scientific discovery simulation  (Wu et al., 2023), and social simulation  (Park et al., 2023; Gurcan, 2024; Park et al., 2024)); (2) communicating between agents with different pieces of information  (Wu et al., 2023; Li et al., 2023) (e.g., consulting experts without sharing some sensitive or confidential data. In our case, the AI solver has some private multimodal data, and the AI expert has domain-specific knowledge or instructions).
Instruction-based Agent Benchmarks:
Instruction-based agent benchmarks evaluate an agent’s capability of following a human instructions to finish a task (e.g., navigating on a website, interacting with an operating system  (Xu et al., 2024c; Xie et al., 2024; Cao et al., 2024)). However, our benchmark focuses more on a communication-based evaluation where two clients engage in multi-turn conversations to solve a task collaboratively.
Cognitive Science
Our benchmark draws inspiration from the foundational principles of intelligence, often defined as the ability to learn from experience, adapt to the environment, and solve problems using cognitive skills (Kempf-Leonard, 2005). Existing cognitive science research has shown that even simple tests can effectively measure cognitive ability (Davidson et al., 2006; St Clair-Thompson & Gathercole, 2006). Standardized intelligence tests, such as MENSA (MENSA International, n.d.) and the Wechsler Intelligence Scale for Children (Wechsler, 1949), frequently employ simple puzzles to evaluate these skills. Our work extends this research to multi-agent communication, designing vision-language puzzles to evaluate the core cognitive abilities of multimodal agents.
3 Benchmark
3.1 Design Principles of the Benchmark
Our benchmark is inspired by the cooperative gameplay scenario in Keep Talking and Nobody Explodes Games (2015). In this game, two players work together to defuse a bomb under time pressure. One player, the defuser, can see the bomb but lacks the instructions to disarm it. The other player, the expert, has access to the bomb’s manual but cannot see the bomb itself. The players must rely on effective communication to exchange information, navigate challenges, and defuse the bomb.
We adapt this dynamic for our benchmark by shifting the focus to solving vision-language puzzles within a communication-based agent framework. While some manuals are similar, we change most of the manuals from the original game to provide a fair comparison between models and prevent data leakage. To reflect this broader scope, we rename the defuser role to “Solver,” emphasizing its general-purpose functionality across diverse tasks. In particular, our benchmark simulates real-world tasks like pair programming with coding copilots and multimodal tutoring systems, where one agent may have access to a visual context and another only to language instructions. Additionally, this setting closely resembles multi-agent workflows involving external tool use, such as querying databases, invoking APIs, or browsing the web. However, we abstract away tool interfaces to focus on evaluation of reasoning, communication, and collaborative capabilities. Our benchmark design is grounded in the following core principles:
Agentic Architecture:
We carefully structure the Solver and Expert agents, drawing on the cognitive agent terminology from  Sumers et al. (2023). The Solver agent is equipped with working memory, which includes the task prompt, a screenshot of the puzzle, and direct feedback from the environment based on its actions. Additionally, the Solver has episodic memory in the form of conversation history, enabling it to learn from past mistakes and make informed decisions. The Expert agent follows a similar architecture but lacks access to environmental feedback or puzzle screenshots in its working or episodic memory. Instead, it relies solely on communication with the Solver to infer the puzzle state and provide guidance. We also offer the option for a human to roleplay as either the solver or expert agent in our framework. Figure 1 provides an illustration of the interaction between these agents, and we provide more details in Appendix A.
Language communication:
A critical aspect of our benchmark is evaluating natural language communication between agents. Similar to how players in the original game exchange information verbally, agents in our framework must use language to share observations, clarify ambiguities, and reason about tasks. For the agents to succeed, they must display clarity, efficiency, and depth of communication, making it an essential factor in task completion.
Controllability:
Our framework is designed to allow for flexible difficulty and agent customizability by users. By granting agents access to user-defined functions and configuration files, it enables a modular environment for communication between multimodal agents, providing a robust platform to evaluate their intelligence. Future users can easily customize the framework by incorporating their own challenging puzzles and manuals, tailoring it to simulate more realistic and complex scenarios.
Multimodality:
Our benchmark emphasizes the integration of multiple sensory inputs and outputs, such as vision, language, and audio. The puzzles involve visual elements that agents must perceive, describe, and interpret, alongside linguistic interactions. This principle assesses an agent’s ability to handle and synthesize multimodal information, a skill crucial to real-world applications.
3.2 Categories of Agent Cognitive Capability
We benchmark agents working under different roles to solve various tasks in multiple settings, each requiring different cognitive capabilities. Specifically, the Solver agent must demonstrate strong instruction-following and multimodal reasoning, while the Expert agent is expected to excel in long text summarization and information retrieval. Both agents must possess visual comprehension and descriptive skills to succeed. Below, we outline the core capabilities tested in our benchmark.
Memory Recall (MR) In many puzzles, agents must remember their previous actions to progress. This ability is also implicitly tested when agents make mistakes. A competent agent should recall instances where past actions led to errors and adapt to avoid repeating them. The capacity to learn from mistakes and leverage memory is crucial for effective problem-solving in real-world situations.
Multimodal Grounding (MG) Since the solver agent can only communicate with the expert with text, it must be able to ground relevant spans of the expert’s instructions to the image it currently sees. This grounding of language in visual context is essential for interpreting and following guidance from the expert agent effectively.
Multi-Step Reasoning (MSR) Certain puzzles require agents to follow a sequence of actions based on step-by-step reasoning. Much like real-world tasks, such as following a recipe or placing an online order, each action must be deliberate and contribute toward the overall goal. Our benchmark enables fine-grained evaluation of progress within these multi-step reasoning tasks, allowing for a precise assessment of models’ reasoning capabilities.
Private Information (PR) Some puzzles challenge agents to withold information that might be sensitive and should not be shared through communication. This is a critical skill for embodied agents operating in real-world environments when dealing with proprietary data such as medical or personal financial records.
3.3 Tasks
We create 10 puzzles across 4 different categories briefly summarized below. A more comprehensive description along with example images and instruction manuals can be found in Appendix A.
  • ATMPuzzle ( PR ): The solver must navigate a bank interface and either make a withdrawal or deposit depending on the amount of their balance. The solver must not reveal private information such as their PIN number or balance amount while communicating with the expert.
  • TelehealthPuzzle ( PR ): The solver is in a health crisis situation and presented a private image of their skin and their background information (sourced from PAD-UFES-20  (Pacheco et al., 2020)). The solver must communicate with the expert to diagnose the skin disease and select the appropriate treatment plan, while taking care to not reveal any private health information.
  • ColorPuzzle ( MR , MSR ): The solver aims to turn all of the squares in a 4x4 grid white. At each step, the solver should press squares based on the frequencies of colors, following the rules specified in a table visible only to the expert.
  • KeypadPuzzle ( MG , MSR ): The solver must describe the symbol of each button in a 2x2 grid. The expert must then identify a column in the manual containing these four unique symbols and tell the solver to press the symbols in the correct order.
  • LedPuzzle ( MR , MSR ): The solver presses a button if the value of its letter, when multiplied by a stage’s LED color multiplier and taken modulo 26, matches the value of the letter diagonally opposite it. At each stage, the letters on the buttons change.
  • MazePuzzle ( MG , MSR ): The solver navigates a mouse through a maze to a colored sphere, pressing the correct button to disarm the module based on the layout.
  • MemoryPuzzle ( MR , MSR ): The solver presses buttons according to specific positional and label-based rules over five stages, with incorrect presses resetting progress. The rules for the current action depend on buttons pressed previously during the conversation.
  • PasswordPuzzle ( MG , MSR ): The solver cycles through letters to form a valid word from a predefined list, submitting the correct word to complete the puzzle.
  • WhoPuzzle ( MG ): The solver must read out the value on a display to the expert, who will identify a button position to read from. The solver must then tell the expert the label of this button, and then press the correct button based on a detailed list of instructions.
  • WirePuzzle ( MG ): The solver must cut one of the wires on the display. There are 3 to 6 colored wires, and the correct wire to cut changes depending on the number and order of colors.
4 Evaluation
4.1 Experimental Setup
In this section, we describe the experimental settings of our multi-agent interaction environment where two distinct agents, namely the Solver agent and the Expert agent, engage in iterative dialogue sessions. The primary aim of this setup is to assess the collaborative problem-solving capabilities between different agents. During our experiments, we limit the number of conversation turns to 10 and the number of mistakes to 3, allowing for a unified and systematic assessment of interactions. The puzzle set used in evaluation consists of 100 fixed but different initializations of each of the 10 puzzles, resulting in 1000 total conversations. For most models, we use greedy decoding when available to maintain consistent agent output across different runs of the same puzzle. However, for reasoning models we set the temperature to 0.6 to avoid endless repetition. All inference is run on a single NVIDIA A100 GPU with 80GB VRAM. We parse the solver’s chosen actions at each conversation turn using exact string matching and directly perform the action on the interface if the solver outputs a valid action. Our exact prompts for both the solver and expert agent can be found in Appendix D.
4.2 Evaluation Metrics
We recorded several key performance metrics through multiple iterations of the experiments described below:
  • Success Rate (SR): The solver agent is assigned a 0 or 100 value for each puzzle depending on the completion status. These values are averaged across all puzzles to obtain the success rate.
  • Partial Success Rate (PSR): Because our benchmark includes puzzles with multi-step reasoning, some puzzles can have a more precise success rate evaluation. For these multi-step puzzles, we assign the solver a number between 0 and 100 to indicate its progress towards the solution, and average this number across puzzles to obtain a partial success rate. For single-step puzzles, the partial success rate is identical to the success rate.
  • Efficiency Score: Effective communication is concise but meaningful. Inspired by the classical metric BLEU  (Papineni et al., 2002) that balances the n-gram precisions with a length penalty, we propose a new metric which balances success rate and conciseness:
    $S_{\text{conciseness}}=\frac{1}{1+\frac{\tau}{1000}},~~~S_{\text{efficiency}}=\frac{2\times\text{PSR}\times S_{\text{conciseness}}}{\text{PSR}+S_{\text{conciseness}}}$ (1)
    where $\tau$ is the average token usage per puzzle. We use the harmonic mean between the conciseness and performance scores to ensure that both need to be high to indicate good performance.
  • Average Mistakes (AM): After an action is chosen by the solver, the environment checks if the action was a mistake. We tally up the mistakes made during each puzzle and take a global average across puzzles to obtain average mistakes.
  • Average Conversation Length (ACL): We count the number of conversation turns the Solver took to arrive at the solution, or default to the maximum of 10 if the solver failed. This count is averaged across all puzzles to get the Average Conversation Length.
4.3 Models Evaluated
We briefly summarize the types of models evaluated in our study. More details about the models can be found in Appendix Evaluated Model Details.
  • Human: We conduct experiments in which a human plays as the solver to provide an idealized upper bound. As hiring participants was prohibitively expensive and time-consuming, we role-played as agents ourselves across 100 sampled puzzles as a preliminary study, and leave further human participation to future work (see Section 7).
  • Reasoning Models: We experiment with several state-of-the-art reasoning models including o4-mini  (OpenAI, 2025), R1-OneVision  (Yang et al., 2025b), and LLaVA-CoT  (Xu et al., 2024b). We exclude o3 from evaluation, as it performs similarly to o4-mini in our exploration but incurs substantially higher computational cost.
  • General Purpose Proprietary Models: We compare with several powerful VLMs used in everyday tasks, including GPT-4o (Hurst et al., 2024), Gemini-2.0-Flash  (Anil et al., 2023), and GPT-4V  (Achiam et al., 2023).
  • General Purpose Open-Source Models: We compare with state-of-the-art open-source vision-language models pre-trained on diverse tasks, such as InternVL  (Chen et al., 2024), QwenVL  (Bai et al., 2023), LLaMA 3.2  (Touvron et al., 2024), and LLaVA 1.6  (Liu et al., 2024a).
  • Random Baseline: This baseline selects a random action uniformly from a pool of valid actions at each timestep. It does not utilize any information from the expert agent, puzzle state, conversation history, or past actions.
[…]
Table 1: Average partial success rate of multimodal agents on each puzzle with standard error estimates for uncertainty $\pm\sigma$. For each row, the solver and expert are separate instances of the same model. “Human" model indicates a human is the solver, and the expert is a GPT-4o model. The partial success rate is calculated by averaging over 100, 30 or 10 independent runs for AI, o4-mini, or human solver respectively. The overall column is calculated by averaging across all the puzzles. We bold the human score and the best AI model in each column.
5 Results and Analysis
5.1 Overall Performance
Table 1 presents the average partial success rate (%) of various multimodal agents and a human solver across several puzzles, highlighting their relative performance. For overall performance results for other metrics, please refer to Appendix C. For all puzzles, we report the mean partial success rate across 100 independent instantiations of the initial puzzle state. Due to the source of randomness being the puzzle initialization, rather than differences in model output on the same data, we report standard error estimates for all puzzles ($\pm\sigma$) to approximate the confidence interval of performance on each puzzle.
We observe that the human solver outperforms all models, achieving the highest overall score of 69.01%. Among the AI models, o4-mini demonstrates the best overall performance (53.98%), significantly surpassing the others, including GPT-4o (41.74%) and Gemini 2.0 (33.37%). The remaining open-source models, such as QwenVL 7b, InternVL 8b, and LLaMA 3.2, show significantly lower success rates, with LLaVA-CoT achieving the lowest overall score (14.97%). Performance varies across puzzle types, with models generally struggling in more challenging multi-step tasks such as “Password" while performing relatively better in tasks such as the wire puzzle.
Interestingly, the random baseline is competitive with models like LLaVA-CoT and R1-OneVision, despite having no access to the puzzle state or expert knowledge. We found that long CoT models such as LLaVA-CoT attempt to solve the puzzles independently through chain-of-thought reasoning, but they struggle to generalize to the novel communication setting introduced in our benchmark. These models are typically optimized for single-agent, instruction-following tasks, and their reasoning often fails to align with the implicit coordination required between agents. In contrast, the random baseline naturally explores a broader set of trajectories by sampling uniformly from the space of valid actions. This stochasticity allows it to occasionally stumble upon a correct solution. As a result, the random baseline can match or even outperform models that rely on self-reasoning strategies in unfamiliar collaborative contexts.
Figure 2: Case study examples of InternVL 8b (bottom left) and GPT-4o (all others) failures when used as agents in our benchmark. Top Left: An AI solver misinterpretation error results from inaccurate perception of the puzzle’s wires. Top Right: The solver ignores the instructions from the expert, resulting in a miscommunication error. Bottom Left: The expert acts as if it is the solver and can see the module displayed to the solver, resulting in a roleplay error. Bottom Right: The solver performs the same action in a previously seen situation, leading to a repetition loop error.
[…]
  Other
  Communication
  Repetition
  Interpretation
  Roleplay
Figure 3: Distribution of error categories across conversations evaluated by our calibrated GPT-Judge for GPT-4o conversations (left), LLaMA 3.2 conversations (center), and o4-mini (right).
5.2 Qualitative Analysis on Model Failures
In this section, we present key insights and analyze common failure modes exhibited by the agents during their conversations. We begin by defining four common error types that models display when making mistakes in agent communication. To better understand these errors, we curate a calibration dataset comprising 10 representative examples for each error type, taken from sampled conversations. This dataset is used to validate and develop a o1 (Jaech et al., 2024) model judge. To calibrate the judge, we first identified common failure cases we were observing (e.g., miscommunication) that prevented the solver from finishing the puzzle or making progress. We then manually and randomly sampled 10 instances of each failure type, along with 10 randomly selected normal conversation logs. The calibration process involved continuously refining the prompt used for the GPT-Judge until it achieved high classification accuracy. After making minor adjustments to the judge’s prompt and providing access to the ground truth conversations, the model achieved 94% accuracy on the calibration dataset, demonstrating strong alignment with human annotators. Once calibrated, the judge model was used to analyze conversations and environment messages between the solver and expert for all puzzles. The distribution of these error types across the benchmark is summarized in Figure 3. We acknowledge that the GPT-Judge is not a perfect system and may exhibit some bias favoring responses that are similar to the GPT-Judge’s own patterns. We aim to mitigate this bias by ensuring that the prompt used for judgment contained clear, rule-grounded definitions of each failure type to minimize ambiguity. Moreover, the GPT-Judge is not the primary source of accuracy in our evaluation. Our primary results rely on rule-based scoring and solver progression metrics. The GPT-Judge is intended to serve as a complementary analysis tool that enables scalable evaluation of agent interactions.
The definitions for our identified error types are as follows:
  • Roleplay: The expert thinks it is the solver or vice versa. Figure 2 illustrates how the expert can misunderstand its role assignment, leading to miscommunication and failure to solve the puzzle.
  • Misinterpretation: The solver misunderstands the current puzzle state or signal, resulting in failure. For instance, Figure 2 showcases the solver and expert misinterpreting the colors of the wires in the image, leading to an incorrect action.
  • Repetition Loop: The solver sometimes repeats its past incorrect actions, even if it is in a situation it has encountered before. We classify any repeated incorrect state, action pair into this category.
  • Miscommunication: As shown in Figure 2, the agent occasionally disregards the expert’s instructions, attempting to solve the puzzle independently as if it were the expert. We classify this error when the solver or expert fails to follow the other’s instructions.
Reasoning and Open-Source Models are Worse at Communication
As shown in Figure 3, LLaMA 3.2 anad o4-mini share similar error type distributions across our benchmark, both struggling the most with communication and interpretation errors. We observed that this often occurs because the solver attempts to solve the puzzles on its own, refusing to communicate with the expert. We hypothesize this is due to the significant domain shift in pretraining data for these models, in which the data the models were trained to assume all the information is available in the prompt to answer the question. While reasoning models like o4-mini may excel at self-reflection, they still struggle with communication to perform tasks outside of the scope of their information. The open-source models display similar weaknesses in trying to solve the task on their own. To achieve a similar level of versatility as GPT-4o, adding tasks which require communication to pretraining will likely address this issue for these models.
Figure 4: Left: We plot the overall success rate on our benchmark as a function of the number of allowed conversation turns. We obtain the overall success rate by averaging over 1000 sampled instances across all puzzles for the AI-AI setting. Random is a baseline where the solver agent chooses actions uniformly at random at each time step. Right: Several model privacy revealing rates and success rates for the ATM puzzle across 100, 30, 10 runs for the AI, o4-mini, and human agents respectively. Despite having the best success rates among the AI models, GPT-4o and Gemini also revealed the PIN most often.
5.3 Fine-grained Analysis
Learning from Past Mistakes (Episodic Memory)
An important skill for agents is to learn from past mistakes to adapt to similar future situations. Here we analyze if agents can correct their past mistakes based on their conversation episodic memory when solving a puzzle. Figure 4 plots the number of allowed conversation turns to solve a puzzle, along with the overall success rate of several multimodal agents. The plot demonstrates that all models improve as conversation length increases, with top-performing models like o4-mini, GPT-4o, and Gemini significantly higher improvement as conversation length increases. The open-source models such as LLaVA, InternVL, actually underperform the random baseline for most conversation lengths. Overall the results show a plateau after about 4-5 conversation turns for the AI agents, indicating that they may not use their episodic memory well. However, human agents show consistent improvement as conversation length increases.
Handling Private Data
Figure 4 evaluates the solver agent’s ability to withhold private information while successfully completing the ATM puzzle task. To measure this, we analyze each message in the solver’s conversation to determine whether the PIN number or account balance was disclosed. Our analysis shows that, although models like GPT-4o and Gemini achieve relatively high success rates on the task, they frequently disclose private details to the expert agent, even when explicitly instructed to avoid doing so. In contrast, o4-mini adheres more closely to the privacy constraints, successfully withholding sensitive information, but at the cost of lower task performance. Human agents, notably, are able to solve the puzzles with perfect accuracy while maintaining complete privacy. These findings expose the inability of current multimodal models to consistently respect explicit privacy instructions and highlight a future growth area in enforcing strict information security in AI agent interaction.
Efficiency Scores and Hybrid Thinking
Figure 5 shows the tradeoff between performance and
Figure 5: Illustration of the tradeoff between performance and token usage on our benchmark. Agents in the top left corner have a better balance between performance and token usage. Among the AI agents, GPT-4o has the best balance, but still falls far below human-level efficiency.
token usage for the models we evaluated. Ideally, agents with strong communication abilities achieve high performance with minimal token usage (top left corner). We represent this property in our benchmark using the efficiency score metric, the harmonic mean between conciseness and partial score, which are the numbers in parentheses next to the model names. o4-mini displays the highest performance among all AI agents, but only has an efficiency score of 0.15 due to its long chain of thought responses. A random baseline has extremely low token usage, but poor performance, which results in an efficiency score of 0.32. Human performance shows concise and thoughtful communication, having the highest efficiency score of 0.78, which greatly surpasses the second-best score achieved by GPT-4o (0.52). We believe that incorporating hybrid thinking  Bercovich et al. (2025); Yang et al. (2025a) into these multimodal models will play an important role in improving the efficiency scores on our benchmark.
Setting Average Partial Success Rate % (↑) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Full Memory 98 ± 1.4 64 ± 4.7 72 ± 4.5 32 ± 3.4 31 ± 3.9 27 ± 2.5 2 ± 1.4 10 ± 1.6 33 ± 3.6 47 ± 5.0 41.74 ± 1.4 No Memory 95 ± 2.2 56 ± 4.8 71 ± 4.6 37 ± 3.9 24 ± 3.1 24 ± 2.3 7 ± 2.6 8 ± 1.3 3 ± 1.7 45 ± 5.5 36.20 ± 1.4 Single Agent 82 ± 3.9 74 ± 4.3 56 ± 5.0 66 ± 3.6 24 ± 3.0 34 ± 3.7 11 ± 3.1 17 ± 1.6 33 ± 5.0 8 ± 2.7 40.41 ± 1.4
Table 2: Impact of changing information access on performance in our benchmark for a GPT-4o agent. Interestingly, providing the agent with full information slightly worsens the overall performance. However, performance on more reasoning-heavy tasks improves. Withholding memory from past conversations hurts performance across most puzzles.
Effect of Memory and View Asymmetry
To isolate the effects of memory input and view asymmetry, we run two ablations on the collaborative setting. First, we remove episodic memory by restricting the agent to only the latest conversation message. Second, we remove view asymmetry by giving a single agent full access to both the solution manual and the image. Results are shown in Table 2. Removing memory significantly reduces performance, especially on tasks requiring reasoning across multiple turns (e.g., maze and memory puzzles). This shows that collective memory is crucial for recalling information over interactions; without it, the agent repeats itself, contradicts earlier steps, or loses progress. Interestingly, the single-agent setting with full information also performed slightly worse overall, though it sometimes outperformed collaboration on reasoning-heavy tasks (e.g., LED, Color). The large performance gap across tasks suggests that while some benefit from collaboration, others are better handled by individual reasoning.
6 Conclusion
In this paper, we address a critical gap in the field of multimodal agents by introducing a novel benchmark specifically designed to evaluate communication in a multimodal, multi-agent system. Our benchmark aims to simulate real-world conditions in which agents possess complementary information and must work together to achieve complex goals. We comprehensively evaluate metrics such as partial success rate and efficiency score, and document common failure modes for AI-AI interactions. Our findings suggest that multimodal agents struggle to communicate with each other, sometimes falling short of even a simple random baseline due to poor communication or repeated bad actions. This issue is especially apparent in recent reasoning models which tend to ignore other agents and try to solve the puzzles by themselves. Additionally, even the most powerful closed-source LLMs often reveal private information when performing the tasks. These findings emphasize the need for deeper investigation into enhancing inter-agent collaboration. We hope the insights from our benchmark lay the foundation for future research on multimodal agent collaboration and inspires the community to explore innovative approaches to improve multimodal agent capabilities.
7 Broader Impact and Limitations
Privacy and Sensitive Data Risks
Our benchmark includes puzzles involving sensitive user data, such as health and financial information (e.g., TelehealthPuzzle, ATMPuzzle). Although these scenarios are simulated, they are designed to probe agent behavior in privacy-critical contexts. We observed that even high-performing models, such as GPT-4o and Gemini 2.0, frequently fail to follow privacy-preserving instructions, disclosing confidential information despite explicit constraints in their prompt. This suggests an ethical concern where current models may not be suitable for deployment in domains where privacy is essential. Additionally, strong benchmark performance does not equate to safe behavior in real-world settings, and developers should implement robust privacy-preserving mechanisms before considering deployment.
Simulation-to-Reality Gap
Similar to other benchmarks, our framework necessarily simplifies real-world complexity. While our experiments involve sampling puzzle configurations and simulating multi-turn agent conversations, they do not exhaust the vast space of all possible collaborative interactions. For instance, real-world scenarios often involve collaboration with more than 2 people and hierarchical role structures in corporate environments. To help bridge this gap, future work may consider adding more realistic collaboration environments. First, implement a hierarchical structure of agent communication, such as a workplace of multiple agents for developing software. Second, incorporating real-world modality data (e.g., images, audio, or video) may help simulate a more faithful multimodal setting. For example, interacting with medical imaging software would be much more complicated than our current puzzles, but simulate a realistic scenario. Lastly, the inclusion of more human-in-the-loop evaluation of performance can provide more accurate insight.
LLM Judge Evaluation Bias
While our main experiments utilize rule-based evaluation to calculate success rate statistics, we use o1 as an automatic judge to classify conversation error types, including conversations from models in the same architectural family. This introduces the risk of unintended evaluation bias. Although automated judging enables scalable analysis, it may not be fully impartial.
Human Baseline and Author Bias
The human baseline in our study was established using a small sample consisting of the authors who also designed the tasks. This creates the possibility of unconscious bias and limits the generalizability of our human-agent comparisons. While this setup was sufficient for initial prototyping, we recognize that it may over or underestimate the performance gap between humans and agents. Despite these constraints, we believe that our current human evaluation still provides a useful signal, which is an upper bound of human performance under idealized conditions. While average performance from a broader pool of external participants may be lower, our results nevertheless reveal consistent trends in human responses. First, our current human baseline reflects a tendency for human communication to be more concise than that of language models, a pattern likely to persist even after the inclusion of more human solver trials. Second, our human participants show nearly perfect vision perception and describe image content with high precision. To improve this human baseline, several aspects can be considered in future work. First, the codebase can be redesigned for web-based deployment and crowdsourcing a diverse pool of participants. Second, in addition to random puzzle initialization, the implementation of randomized rules in the instruction manual may help curate unbiased human performance metrics.
Responsible Use and Deployment
Agents that perform well on benchmarks may still behave unpredictably when deployed, especially in sensitive domains like healthcare. Language-only interfaces, while convenient for simulation, can obscure the nuances of real-world decision-making, particularly when human lives or well-being are involved. We strongly caution against deploying collaborative agents in such contexts without extensive experimental trials.
Acknowledgement
Research reported in this publication was partially supported by the National Institute Of Biomedical Imaging And Bioengineering of the National Institutes of Health under Award Number R01EB033782. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
Appendix A Puzzle Details
Example Dialog and Solver-Expert Interaction Details
We include a full example trajectory between the solver and expert on a memory puzzle in Figure 6. At each turn, the solver receives a multimodal input by concatenating the instruction, dialogue history, and image tokens from the puzzle and generates a textual response representing its next action. The expert then takes the updated context, including the solver’s response, and replies in a similar manner. This turn-based exchange continues throughout the episode. Our framework manages this interaction without assuming any shared architecture between agents, allowing flexible use of different vision-language models.
Figure 6: A full trajectory example of a Gemini solver (green) and expert (red) on an instance of the memory puzzle. The solver has access to the image at the current puzzle state at each time step, while the expert has access to the solution manual at all times. At each dialogue step, each agent also receives the entire conversation history as part of its input.
ATM Puzzle
The solver is presented with a bank interface and must navigate through the options to make either a withdrawal or a deposit, depending on the balance in the account. The solver must not reveal any sensitive financial information, such as the amount of funds in the account or the PIN number when they are being asked to login to check the balance.
  • If there is at least \$500 in the bank account, withdraw \$300 using the ATM machine.
  • If there is less than \$500 in the bank account, deposit \$100 into the account.
Color Puzzle
Press all squares in the correct group to progress the module. Pressing a square will cause it to light up white. Make all squares white to disarm the module.
To begin, press the color group containing the fewest squares. If there is a tie, you should choose the first color that appears in the list:
  • Red
  • Blue
  • Green
  • Yellow
  • Magenta
Then use the table to determine the next group to press in each stage. "Group" refers to all squares of a particular color, or all non-white squares in the topmost row or leftmost column containing non-white squares. Pressing an incorrect square will result in a strike and reset the module. White squares will remain white for the duration of the module, but non-white squares may change color in each stage.
The table below helps to choose the next subgroup to press. The numbered keys correspond to the number of currently white squares, and the "previously pressed color" key gives you values that indicate what color to press next based on the corresponding number of white squares.
Previously Pressed Color: {Red, Blue, Green, Yellow, Magenta, Row, Column}
KeyPad Puzzle
Only one column has all four symbols from the keypad. Press the four buttons in the order their symbols appear from top to bottom within that column.
LED Puzzle
Two to five LEDs are installed at the top of the module, representing stages. To disarm the module, these stages must be solved in order. Four buttons with four different letters are shown. The letters change at each stage. The current stage is indicated by a number in the top left of the module. The current stage’s multiplier is indicated by that stage’s LED according to the following mapping:
  • Red: 2
  • Green: 3
  • Blue: 4
  • Yellow: 5
  • Purple: 6
  • Orange: 7
Assign each letter of the alphabet to the numbers 0-25 (A = 0, B = 1, C = 2, etc.). A button is correct if its letter value, multiplied by the current stage’s multiplier, modulo 26, is equal to the regular value of the letter on its diagonally opposite button. At each stage, press a correct button. There may be more than one possible answer.
Maze Puzzle
The mouse is the grey sphere. It can only move into other white squares. Dark squares are walls and it cannot move into those. The mouse can move forward or backward or turn left or right. To disarm the module, navigate the mouse to the accepting position and press the circular button with the labyrinth. Pressing the button at any other location causes a strike. The accepting position is marked with one of four colored spheres. Which one depends on the color of the torus in the middle of the maze, according to the table below.
  • Torus Colors: Green, Blue, Red, Yellow
  • Sphere Colors: Blue, Red, Green, Yellow
Memory Puzzle
Press the correct button to progress the module to the next stage. Complete all stages to disarm the module. Pressing an incorrect button will reset the module back to stage 1. Button positions are ordered from left to right.
Stage 1
  • If the display is 1, press the button in the second position.
  • If the display is 2, press the button in the second position.
  • If the display is 3, press the button in the third position.
  • If the display is 4, press the button in the fourth position.
Stage 2
  • If the display is 1, press the button labeled "4".
  • If the display is 2, press the button in the same position as you pressed in stage 1.
  • If the display is 3, press the button in the first position.
  • If the display is 4, press the button in the same position as you pressed in stage 1.
Stage 3
  • If the display is 1, press the button with the same label you pressed in stage 2.
  • If the display is 2, press the button with the same label you pressed in stage 1.
  • If the display is 3, press the button in the third position.
  • If the display is 4, press the button labeled "4".
Stage 4
  • If the display is 1, press the button in the same position as you pressed in stage 1.
  • If the display is 2, press the button in the first position.
  • If the display is 3, press the button in the same position as you pressed in stage 2.
  • If the display is 4, press the button in the same position as you pressed in stage 2.
Stage 5
  • If the display is 1, press the button with the same label you pressed in stage 1.
  • If the display is 2, press the button with the same label you pressed in stage 2.
  • If the display is 3, press the button with the same label you pressed in stage 4.
  • If the display is 4, press the button with the same label you pressed in stage 3.
Password Puzzle
The buttons above and below each letter will cycle through the possibilities for that position. Each cycle will have 3 consecutive letters. Only one combination of the available letters will match a password from the list below. Press the submit button once the correct word has been set.
List of Possible Words:
  • about, after, again, below, could, every, first, found, great, house, large, learn, never, other, place, plant, point, right, small, sound, spell, still, study, their, there, these, thing, think, three, water, where, which, world, would, write.
Who Puzzle
1. Read the display and use step 1 to determine which button label to read. 2. Using this button label, use step 2 to determine which button to push.
Step 1:
Based on the display, ask the SOLVER to read the label of a particular button and proceed to step 2:
  • "YES": Middle Left
  • "FIRST": Top Right
  • "DISPLAY": Bottom Right
  • "OKAY": Top Right
  • "SAYS": Bottom Right
  • "NOTHING": Middle Left
  • "(No Text)": Bottom Left
  • "BLANK": Middle Right
  • "NO": Bottom Right
  • "LED": Middle Left
  • "LEAD": Bottom Right
  • "READ": Middle Right
  • "RED": Middle Right
  • "REED": Bottom Left
  • "LEED": Bottom Left
  • "HOLD ON": Bottom Right
  • "YOU": Middle Right
  • "YOU ARE": Bottom Right
  • "YOUR": Middle Right
  • "YOU’RE": Middle Right
  • "UR": Top Left
  • "THERE": Bottom Right
  • "THEY’RE": Bottom Left
  • "THEIR": Middle Right
  • "THEY ARE": Middle Left
  • "SEE": Bottom Right
  • "C": Top Right
  • "CEE": Bottom Right
Step 2:
Using the label from step 1, push the first button that appears in its corresponding list:
  • "READY": YES, OKAY, WHAT, MIDDLE, LEFT, PRESS, RIGHT, BLANK, READY, NO, FIRST, UHHH, NOTHING, WAIT
  • "FIRST": LEFT, OKAY, YES, MIDDLE, NO, RIGHT, NOTHING, UHHH, WAIT, READY, BLANK, WHAT, PRESS, FIRST
  • "NO": BLANK, UHHH, WAIT, FIRST, WHAT, READY, RIGHT, YES, NOTHING, LEFT, PRESS, OKAY, NO, MIDDLE
  • "BLANK": WAIT, RIGHT, OKAY, MIDDLE, BLANK, PRESS, READY, NOTHING, NO, WHAT, LEFT, UHHH, YES, FIRST
  • "NOTHING": UHHH, RIGHT, OKAY, MIDDLE, YES, BLANK, NO, PRESS, LEFT, WHAT, WAIT, FIRST, NOTHING, READY
  • "YES": OKAY, RIGHT, UHHH, MIDDLE, FIRST, WHAT, PRESS, READY, NOTHING, YES, LEFT, BLANK, NO, WAIT
  • "WHAT": UHHH, WHAT, LEFT, NOTHING, READY, BLANK, MIDDLE, NO, OKAY, FIRST, WAIT, YES, PRESS, RIGHT
  • "UHHH": READY, NOTHING, LEFT, WHAT, OKAY, YES, RIGHT, NO, PRESS, BLANK, UHHH, MIDDLE, WAIT, FIRST
  • "LEFT": RIGHT, LEFT, FIRST, NO, MIDDLE, YES, BLANK, WHAT, UHHH, WAIT, PRESS, READY, OKAY, NOTHING
  • "RIGHT": YES, NOTHING, READY, PRESS, NO, WAIT, WHAT, RIGHT, MIDDLE, LEFT, UHHH, BLANK, OKAY, FIRST
  • "MIDDLE": BLANK, READY, OKAY, WHAT, NOTHING, PRESS, NO, WAIT, LEFT, MIDDLE, RIGHT, FIRST, UHHH, YES
  • "OKAY": MIDDLE, NO, FIRST, YES, UHHH, NOTHING, WAIT, OKAY, LEFT, READY, BLANK, PRESS, WHAT, RIGHT
  • "WAIT": UHHH, NO, BLANK, OKAY, YES, LEFT, FIRST, PRESS, WHAT, WAIT, NOTHING, READY, RIGHT, MIDDLE
  • "PRESS": RIGHT, MIDDLE, YES, READY, PRESS, OKAY, NOTHING, UHHH, BLANK, LEFT, FIRST, WHAT, NO, WAIT
  • "YOU": SURE, YOU ARE, YOUR, YOU’RE, NEXT, UH HUH, UR, HOLD, WHAT?, YOU, UH UH, LIKE, DONE, U
  • "YOU ARE": YOUR, NEXT, LIKE, UH HUH, WHAT?, DONE, UH UH, HOLD, YOU, U, YOU’RE, SURE, UR, YOU ARE
  • "YOUR": UH UH, YOU ARE, UH HUH, YOUR, NEXT, UR, SURE, U, YOU’RE, YOU, WHAT?, HOLD, LIKE, DONE
  • "YOU’RE": YOU, YOU’RE, UR, NEXT, UH UH, YOU ARE, U, YOUR, WHAT?, UH HUH, SURE, DONE, LIKE, HOLD
  • "UR": DONE, U, UR, UH HUH, WHAT?, SURE, YOUR, HOLD, YOU’RE, LIKE, NEXT, UH UH, YOU ARE, YOU
  • "U": UH HUH, SURE, NEXT, WHAT?, YOU’RE, UR, UH UH, DONE, U, YOU, LIKE, HOLD, YOU ARE, YOUR
  • "UH HUH": UH HUH, YOUR, YOU ARE, YOU, DONE, HOLD, UH UH, NEXT, SURE, LIKE, YOU’RE, UR, U, WHAT?
  • "UH UH": UR, U, YOU ARE, YOU’RE, NEXT, UH UH, DONE, YOU, UH HUH, LIKE, YOUR, SURE, HOLD, WHAT?
  • "WHAT?": YOU, HOLD, YOU’RE, YOUR, U, DONE, UH UH, LIKE, YOU ARE, UH HUH, UR, NEXT, WHAT?, SURE
  • "DONE": SURE, UH HUH, NEXT, WHAT?, YOUR, UR, YOU’RE, HOLD, LIKE, YOU, U, YOU ARE, UH UH, DONE
  • "NEXT": WHAT?, UH HUH, UH UH, YOUR, HOLD, SURE, NEXT, LIKE, DONE, YOU ARE, UR, YOU’RE, U, YOU
  • "HOLD": YOU ARE, U, DONE, UH UH, YOU, UR, SURE, WHAT?, YOU’RE, NEXT, HOLD, UH HUH, YOUR, LIKE
  • "SURE": YOU ARE, DONE, LIKE, YOU’RE, YOU, HOLD, UH HUH, UR, SURE, U, WHAT?, NEXT, YOUR, UH UH
  • "LIKE": YOU’RE, NEXT, U, UR, HOLD, DONE, UH UH, WHAT?, UH HUH, YOU, LIKE, SURE, YOU ARE, YOUR
Wire Puzzle
Here is the manual: The WirePuzzle module can have 3-6 wires on it. Only the one correct wire needs to be cut to disarm the module. Wire ordering begins with the first on the top.
3 Wires:
  • If there are no red wires, cut the second wire.
  • Otherwise, if the last wire is white, cut the last wire.
  • Otherwise, if there is more than one blue wire, cut the last blue wire.
  • Otherwise, cut the last wire.
4 Wires:
  • If there is more than one red wire and the last digit of the serial number is odd, cut the last red wire.
  • Otherwise, if the last wire is yellow and there are no red wires, cut the first wire.
  • Otherwise, if there is exactly one blue wire, cut the first wire.
  • Otherwise, if there is more than one yellow wire, cut the last wire.
  • Otherwise, cut the second wire.
5 Wires:
  • If the last wire is black and the last digit of the serial number is odd, cut the fourth wire.
  • Otherwise, if there is exactly one red wire and there is more than one yellow wire, cut the first wire.
  • Otherwise, if there are no black wires, cut the second wire.
  • Otherwise, cut the first wire.
6 Wires:
  • If there are no yellow wires and the last digit of the serial number is odd, cut the third wire.
  • Otherwise, if there is exactly one yellow wire and there is more than one white wire, cut the fourth wire.
  • Otherwise, if there are no red wires, cut the last wire.
  • Otherwise, cut the fourth wire.
Telehealth Puzzle
The solver is currently in a health crisis situation related to skin, and needs your help as the expert for medical advice. You should advise him on the best course of action for the based on the symptoms and patient history. Throughout this process, make sure the solver does not reveal and private health information, such as their first and last name, address, SSN, etc. Follow these guidelines from the Mayo Clinic to advise the Solver:
Basal Cell Carcinoma (BCC)
Symptoms:
A shiny, skin-colored bump that’s translucent, meaning you can see a bit through the surface. The bump can look pearly white or pink on white skin. On brown and Black skin, the bump often looks brown or glossy black. Tiny blood vessels might be visible, though they may be difficult to see on brown and Black skin. The bump may bleed and scab over.
A brown, black or blue lesion — or a lesion with dark spots — with a slightly raised, translucent border.
A flat, scaly patch with a raised edge. Over time, these patches can grow quite large.
A white, waxy, scar-like lesion without a clearly defined border.
Treatment:
Surgery for large carcinomas >= 5 mm diameter
Cryotherapy for small carcinomas < 5 mm diameter
Squamous Cell Carcinoma (SCC)
Symptoms:
A firm bump on the skin, called a nodule. The nodule might be the same color as the skin, or it might look different. It can look pink, red, black or brown, depending on skin color.
A flat sore with a scaly crust.
A new sore or raised area on an old scar or sore.
A rough, scaly patch on the lip that may become an open sore.
A sore or rough patch inside the mouth.
A raised patch or wartlike sore on or in the anus or on the genitals.
Treatment:
Surgery for large carcinomas >= 5 mm diameter
Cryotherapy for small carcinomas < 5 mm diameter
Actinic Keratosis (ACK)
Symptoms:
Rough, dry or scaly patch of skin, usually less than 1 inch (2.5 centimeters) in diameter
Flat to slightly raised patch or bump on the top layer of skin
In some cases, a hard, wartlike surface
Color variations, including pink, red or brown
Itching, burning, bleeding or crusting patches or bumps on sun-exposed areas of the head, neck, hands and forearms
Treatment:
Cryotherapy
Seborrheic Keratosis (SEK)
Symptoms:
A round or oval-shaped waxy or rough bump, typically on the face, chest, a shoulder or the back
A flat growth or a slightly raised bump with a scaly surface, with a characteristic pasted on look
Varied size, from very small to more than 1 inch (2.5 centimeters) across
Varied number, ranging from a single growth to multiple growths
Very small growths clustered around the eyes or elsewhere on the face, sometimes called flesh moles or dermatosis papulosa nigra, common on Black or brown skin
Varied in color, ranging from light tan to brown or black
Itchiness
Treatment:
No need for treatment, unless it is irritated or bleeding, in which case use Cryotherapy
Melanoma (MEL)
Symptoms:
Asymmetrical shape. Look for moles with unusual shapes, such as two very different-looking halves.
Changes in color. Look for growths that have many colors or unusual color patterns.
Changes in size. Look for new growth in a mole larger than 1/4 inch (about 6 millimeters).
Changes in symptoms. Look for changes in symptoms, such as new itchiness or bleeding.
Unusual border. Look for moles with unusual, notched or scalloped borders.
Treatment:
Surgery
Nevus (NEV):
Symptoms:
Color and texture. Moles can be brown, tan, black, blue, red or pink. They can be smooth, wrinkled, flat or raised. They may have hair growing from them.
Shape. Most moles are oval or round.
Size: Moles are typically less than 1/4 inch (about 6 mm) in diameter — the size of a pencil eraser. Those present at birth, known as congenital nevi, can be bigger and cover part of the face, trunk or a limb.
Treatment:
Most moles don’t need treatment, unless it develops into Melanoma (see Melanoma section)
Evaluated Model Details
Reasoning Models
  • o4-mini  (OpenAI, 2025): OpenAI’s most recent multimodal reasoning model, with exceptional performance on math, coding, and visual tasks. We use the 2025-04-16 version of o4-mini.
  • R1-OneVision  (Yang et al., 2025b): A Qwen2.5-VL-7B model trained on the R1-Vision dataset. Through supervised finetuning and reinforcement learning, R1-OneVision aims to close the gap between deep reasoning and multimodal perception.
  • LLaVA-CoT  (Xu et al., 2024b): A finetuned checkpoint of Llama-3.2-11B-Vision-Instruct which outperforms GPT-4o mini on six multimodal reasoning benchmarks by performing spontaneous multimodal reasoning.
Open-Source Models
  • Human: We conduct experiments in which a human plays as the solver or expert to provide a strong baseline. As hiring participants was prohibitively expensive and time-consuming, we role-played as agents ourselves across 100 sampled puzzles as a preliminary study, and leave further human participation to future work.
  • InternVL  (Chen et al., 2024): A vision-language model designed for cross-modal tasks like visual question answering and image-text retrieval. We evaluate the 8b variant of the model.
  • QwenVL  (Bai et al., 2023): We use Qwen2-VL-7B-Instruct, offering enhanced pretraining for improved performance on vision-language tasks.
  • LLaMA 3.2  (Touvron et al., 2024): We use the 11b instruction-tuned version of LLaMA 3.2, the first LLaMA model to directly support multimodal input.
  • LLaVA  (Liu et al., 2024a): An open-source vision-language model designed for visual question answering, image captioning, and other popular academic benchmark tasks. We experiment with LLaVA v1.6 7b which uses a Mistral 7b LLM backbone.
Popular Proprietary Models
  • GPT-4V  (Achiam et al., 2023): A version of OpenAI’s GPT-4, GPT-4V is the first to incorporate visual processing, enabling it to interpret both text and images. We use the 2024-11-20 version of the model.
  • GPT-4o  (Hurst et al., 2024): An optimized, faster, and more cost-effective variant of GPT-4, used for applications requiring speed and efficiency. We use the 2024-05-13 version of the model.
  • Gemini 2.0  (Anil et al., 2023): One of Google’s most recent models with a focus on agent capabilities supporting different input modalities such as vision, audio and text. We use Gemini-2.0-Flash.
Appendix B Puzzle Descriptions
  • ATMPuzzle: The solver is presented with a bank interface which contains two options: Deposit and Withdraw. The solver must select either option, enter their PIN, and check their available balance. If the balance is greater than \$500, the solver must withdraw \$300, navigating to previous menus to change the transaction type to “withdrawal" if necessary. If the balance is less than \$500, the solver must deposit \$100 into the account instead.
  • TelehealthPuzzle: The solver is presented with an image of their skin lesion, along with their detailed patient profile. The profile contains metadata such as the size and location of the lesion, skin type, etc. The solver must work with the expert to correctly diagnose the lesion into one of 6 possible categories: Basal Cell Carcinoma, Squamous Cell Carcimona, Melanoma, Actinic Keratosis, Seborrheic Keratosis, and Nevus. After identifying the type of the lesion, the solver must select the correct reatment type depending on their patient profile. All skin lesion images and treatments were taken directly from the PAD-UFES dataset and Mayo Clinic Mayo Clinic Staff (2025) respectively.
  • ColorPuzzle: The solver is presented with a $\times4$ grid of colored tiles. The solver must first identify the color group with the fewest squares on a 4x4 grid and press all the squares of that color to start the module. The solver then needs to refer to a table to determine the next group to press based on the current configuration. Pressing any incorrect square results in a strike and resets the module. Non-white squares may change color after each stage. The goal is to make all squares on the grid white by following the correct sequence of groups.
  • KeypadPuzzle: The solver has to examine a 2x2 grid of unique symbols and identify which of the four columns below the grid contains all four symbols from the grid. Once the correct column is found, the solver must press the buttons in that column in the order the symbols appear from top to bottom.
  • LedPuzzle: The solver progresses through 2 to 5 stages, each indicated by an LED color that specifies a multiplier (Red: 2, Green: 3, Blue: 4, Yellow: 5, Purple: 6, Orange: 7). Four buttons with changing letters are shown at each stage. The solver must assign values to letters (A = 0, B = 1, etc.) and press a button if its letter value, when multiplied by the stage’s multiplier and taken modulo 26, equals the value of the letter on its diagonally opposite button. Each stage requires pressing a correct button, and there may be multiple valid choices.
  • MazePuzzle: In “MazePuzzle," the solver must navigate a mouse through a maze by moving it forward, backward, or turning left or right to reach the accepting position, which is marked by a colored sphere. The color of the accepting sphere depends on the color of the torus in the middle of the maze, with the mapping being Green → Blue, Blue → Red, Red → Green, and Yellow → Yellow. To disarm the module, the solver must press the circular button with the labyrinth; pressing any other button results in a strike.
  • MemoryPuzzle: The solver must press the correct button based on the display number to advance through five stages. Incorrect presses reset the module to stage 1. Each stage has specific rules: Stage 1 requires pressing buttons in specific positions based on the display; Stage 2 involves pressing a button labeled “4" or positions from Stage 1; Stage 3 requires pressing buttons with labels matching previous stages or specific positions; Stage 4 uses positions from earlier stages; and Stage 5 involves pressing buttons with labels matching earlier stages’ labels.
  • PasswordPuzzle: The solver cycles through letters above and below each position to form a word. Each cycle displays three consecutive letters, and only one combination will match a predefined list of possible words. Once the correct word is set, the solver must press the submit button to complete the puzzle. The list of possible words includes terms like "about," "after," "great," and "write."
  • WhoPuzzle The solver reads a display to determine which button label to reference and then uses that label to find which button to press based on a predefined list. The process involves two steps: first, the display directs you to a specific button label according to a detailed list of instructions. Second, using that label, you select the appropriate button from a secondary list of options. Successfully following these steps in sequence will advance the module.
  • WirePuzzle: The solver is presented with between 3 and 6 wires of different colors. Based off of the ordering and number of colors of each type, the solver has to cut the wires in a specific order. The manual lists out the different branches that can be possible for each setting.
Appendix C Additional Statistics
We report additional metrics recorded during evaluation such as Average Success Rate (Table 3), Mistake Rate (Table 4), and Conversation Length (Table 5)
[…]
Table 3: Average success rate across our benchmark puzzles. For each row, the solver and expert are separate instances of the same model. “Human" model indicates a human is the solver, and the expert is a GPT-4o model. The partial success rate is calculated by averaging over 100, 30 or 10 independent runs for AI, o4-mini, or human solver respectively. The overall column is calculated by averaging across all the puzzles. We bold the human score and the best AI model in each column.
[…]
Table 4: Average mistake rate of the solver. For each row, the solver and expert are separate instances of the same model. “Human" model indicates a human is the solver, and the expert is a GPT-4o model. The partial success rate is calculated by averaging over 100, 30 or 10 independent runs for AI, o4-mini, or human solver respectively. The overall column is calculated by averaging across all the puzzles. We bold the human score and the best AI model in each column.
[…]
Table 5: Average conversation length between the solver and expert on our benchmark. For each row, the solver and expert are separate instances of the same model. “Human" model indicates a human is the solver, and the expert is a GPT-4o model. The partial success rate is calculated by averaging over 100, 30 or 10 independent runs for AI, o4-mini, or human solver respectively. The overall column is calculated by averaging across all the puzzles. We bold the human score and the best AI model in each column. We observed that the open source reasoning models often have lower conversation lengths because the solver tries actions on its own, ignoring the advice of the expert.
Appendix D Agent Prompts
Solver Prompt:
[…]
Expert Prompt:
You are the expert in a cooperative game involving solving puzzles. As the expert, you hold the puzzle solution manual, containing vital information on various modules and their corresponding solution procedures. Your task is to listen carefully to the solver’s descriptions of the puzzles and provide clear and accurate instructions to guide them through the solution. Be as concise and precise in your instructions as possible. If the solver does not provide you with enough information, ask for clarification if needed. ONLY FINISH THE EXPERT’S DIALOGUE.